### v1

# Xuezhang Chen  and Shihong Zhang

Abstract: We introduce the notion of a biharmonic Poisson kernel associated with certain pair of conformal boundary operators and present its explicit formula. With this powerful tool, we next establish the classification theorems of nonnegative solutions to a biharmonic equation on the upper half-space and unit ball with proper pairs of conformal boundary operators.
$\mathbf{2020}$ MSC: 35B53,53C18 (35G30,35J08)
Keywords: Conformal boundary operators, biharmonic Poisson kernel, biharmonic Green function.
1 Introduction
Caffarelli-Gidas-Spruck [5] established the celebrated Liouville-type theorem: For $n\geq 2$, a $C^{2}$ solution of
$-\Delta u=(n+1)(n-1)u^{\frac{n+3}{n-1}},~{}~{}u>0\qquad\mathrm{~{}~{}in~{}~{}}\mathbb{R}^{n+1}$
is of the form
$u(X)=\left(\frac{\varepsilon}{\varepsilon^{2}+|X-X_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\varepsilon\in\mathbb{R}_{+},X_{0}\in\mathbb{R}^{n+1}.$
Under an additional hypothesis $u(X)=O(|X|^{1-n})$ as $|X|\to\infty$, the same result was also proved earlier by Obata [28] and Gidas-Ni-Nirenberg [16]. Let $\mathbb{R}_{+}^{n+1}=\{X=(x,t)|~{}x\in\mathbb{R}^{n},t\in\mathbb{R}_{+}\}$ denote the upper half-space with boundary $\partial\mathbb{R}_{+}^{n+1}=\mathbb{R}^{n}$ and $\mathbb{B}^{n+1}$ denote the unit ball in $\mathbb{R}^{n+1}$. Escobar [13] used the Obata-type argument to characterize minimizers of a sharp Sobolev trace inequality on $\mathbb{B}^{n+1}$, via a conformal map between $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$, which is equivalent to seeking smooth solutions to
$\begin{cases}\displaystyle\Delta u=0,\quad u>0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle-\partial_{t}u(x,0)=(n-1)u^{\frac{n+1}{n-1}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (1.1)
under an extra hypothesis that $u(X)=O(|X|^{1-n})$ as $|X|\to\infty$. Y. Y. Li-Zhu [26] (see also, Y. Y. Li-Zhang [25]) employed the moving sphere method to prove the following Liouville-type theorem without the above hypothesis: A $C^{2}$ solution of (1.1) is of the form
$u(x,t)=\left(\frac{\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}.$
Such Liouville-type theorems of general critical growth second order semilinear elliptic equations on $\mathbb{R}_{+}^{n+1}$ are referred to Escobar [14], Y. Y. Li-Zhang [25] etc.
Sun-Xiong [30] classified nonnegative smooth solutions of the following conformally invariant polyharmonic boundary value problem: For $m\geq 2,n+1>2m$ and $1<p\leq\frac{n+2m-1}{n+1-2m}$,
$\begin{cases}\displaystyle\Delta^{m}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}\Delta^{k}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle(-1)^{m}\partial_{t}\Delta^{m-1}u(x,0)=u^{p}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (1.2)
where $20\leq k\leq m-$ 2. When $m$ is even, assume in addition that $u(X)=o(|X|^{2m-1})$ as $|X|\to\infty$, then for the critical exponent $p=\frac{n+2m-1}{n-(2m-1)}$,
$u(x,t)=$ $c(n,m)\int_{\mathbb{R}^{n}}\frac{t^{2m-1}}{(|x-y|^{2}+t^{2})^{\frac{n+2m-1}{2}}}\left(\frac{\lambda}{\lambda^{2}+|y|^{2}}\right)^{\frac{n-2m+1}{2}}\mathrm{d}y+\sum_{k=1}^{m-1}t^{2k}P_{2k}(x),$
where $x_{0}\in\mathbb{R}^{n},\lambda\in\mathbb{R}_{+}$, $c(n,m)$ is a normalized constant and $P_{2k}(x)$ denotes a polynomial of degree $2k$ in the variable $x$; for $1<p<\frac{n+2m-1}{n+1-2m}$,
$u(x,t)=\sum_{k=1}^{m-1}t^{2k}P_{2k}(x).$
A key step of their approach is to improve the integrability of a Kevin transformation $u^{\ast}$ for $(-\Delta)^{m}$ via an iteration method, and eventually to show that the singularity of $u^{\ast}$ can be removed at $t=0$; in the next step, the integral equation of $u^{\ast}$ on $\partial\mathbb{R}_{+}^{n+1}$ is solved by the classification theorems due to Chen-Li-Ou [9] and Y. Y. Li [24], then the standard method can give the classification of nonnegative solutions to (1.2).
Conformal boundary operators $\mathscr{B}_{k}^{3}$ for $k\in\{0,1,2,3\}$ associated with the Paneitz operator have attracted a lot of attention, see [18, 4, 6, 11] etc. In the study of zeta functional determinants on a four-manifold with boundary, Chang-Qing [8] discovered a third order conformally covariant boundary operator $\mathscr{B}_{3}^{3}$. The discovery of conformal boundary operators $\mathscr{B}_{k}^{3}$ to an $(n+1)$-manifold $(X^{n+1},g)$ with boundary $\partial X^{n+1}=M^{n}$ is due to Branson-Gover [3], Juhl [23], Gover-Peterson [19], Grant [20], Stafford [29] etc. The Paneitz operator $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ enjoy the property (cf. [7, Theorem 1.1]) that if $\tilde{g}=e^{2\tau}g$, then for all $\Psi\in C^{\infty}(\overline{X^{n+1}})$,
$\tilde{P}_{4}(\Psi)=e^{-\frac{n+5}{2}\tau}P_{4}^{g}(e^{\frac{n-3}{2}\tau}\Psi)$ (1.3)
and
~$\tilde{\mathscr{B}}_{k}^{3}(\Psi)=e^{-\frac{n+2k-3}{2}\tau}\mathscr{B}_{k}^{3}(e^{\frac{n-3}{2}\tau}\Psi).$ (1.4)
Readers are referred to Case [7] or authors [11] for explicit formulae of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ on manifolds with boundary. We are concerned with two conformally equivalent model spaces: $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$. On the upper half-space (+$(\mathbb{R}^{n+1}_{+},\partial\mathbb{R}^{n+1}_{+},|\mathrm{d}X|^{2})$, the explicit formulae of conformal boundary operators are $\mathscr{B}^{3}_{0}(u)=u$ and
$\mathscr{B}^{3}_{1}(u)=$ $-\partial_{t}u;$
$\mathscr{B}^{3}_{2}(u)=$ $\partial_{t}^{2}u-\overline{\Delta}u;$
$\mathscr{B}^{3}_{3}(u)=$ $\partial_{t}\Delta u+2\overline{\Delta}\partial_{t}u$
for all $u\in C^{\infty}(\overline{\mathbb{R}_{+}^{n+1}})$. On the unit ball $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, these conformal boundary operators are $\mathscr{B}^{3}_{0}(U)=U$ and
$\mathscr{B}^{3}_{1}(U)=$ $\frac{\partial U}{\partial r}+\frac{n-3}{2}U,$
$\mathscr{B}^{3}_{2}(U)=$ $\frac{\partial^{2}U}{\partial r^{2}}-\Delta_{\mathbb{S}^{n}}U+(n-2)\frac{\partial U}{\partial r}+\frac{(n-3)(n-1)}{2}U,$
$\mathscr{B}^{3}_{3}(U)=$ $-\frac{\partial\Delta U}{\partial r}-2\Delta_{\mathbb{S}^{n}}\frac{\partial U}{\partial r}-\frac{n-3}{2}\frac{\partial^{2}U}{\partial r^{2}}-\frac{3n-5}{2}\Delta_{\mathbb{S}^{n}}U$
$+\frac{n-3}{2}\frac{\partial U}{\partial r}+\frac{(n^{2}-1)(n-3)}{4}U$
for all $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$, where $r=|\xi|$ for $\xi\in\overline{\mathbb{B}^{n+1}}$; the corresponding boundary curvatures $(T_{i}^{3})_{g_{\mathbb{S}^{n}}}$ are constant and set $\mathbb{T}^{3}_{i}=\frac{n-3}{2}(T_{i}^{3})_{g_{\mathbb{S}^{n}}}$ for $1\leq i\leq 3$, explicitly,
$\mathbb{T}^{3}_{1}=\frac{n-3}{2},\qquad\mathbb{T}^{3}_{2}=\frac{(n-1)(n-3)}{2}\qquad\mathrm{and}\qquad\mathbb{T}^{3}_{3}=\frac{(n^{2}-1)(n-3)}{4}$
will be used throughout the paper.
Convention..
Let $\Delta$ and $\overline{\Delta}$ denote the Euclidean Laplacians on $\mathbb{R}^{n+1}$ and $\mathbb{R}^{n}$, respectively. Let $p_{k}^{\ast}=\frac{n+2k-3}{n-3}$ for $k\in\{0,1,2,3\}$, as $p_{k}^{\ast}+1$ is the critical Sobolev exponent associated to $\mathscr{B}^{3}_{k}$.
This motivates us to classify nonnegative smooth solutions to biharmonic boundary value problems on $\mathbb{R}_{+}^{n+1}$ associated with pairs of conformal boundary operators $(\mathscr{B}^{3}_{i},\mathscr{B}^{3}_{j})$ for $i<j\in\{1,2,3\}$: For $p_{i},p_{j}>0$, provided $u(x,0)$ has some suitable decay at infinity,
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(u)=\mathbb{T}^{3}_{i}u^{p_{i}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(u)=\mathbb{T}^{3}_{j}u^{p_{j}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.5)
The nonlinear interaction among $\mathscr{B}^{3}_{i}$ and $\mathscr{B}^{3}_{j}$ brings us new obstacles, compared with vanishing lower order part of derivatives in [30]. So, it has great difficulty in applying the approach in [30] to (1.5), as well as some peculiar phenomena occur.
An extremely powerful tool in our approach is so-called biharmonic Poisson kernel consisting of two kernel functions corresponding to a proper pair of conformal boundary operators $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$. In the following, we separate our discussion into two cases: $n\neq 3$ and $n=3$. In the latter case, it involves the critical dimension for both the Paneitz operator and $\mathscr{B}_{3}^{3}$.
Definition 1.1 (Poisson kernel for a biharmonic boundary value problem).
For $i,j\in\{0,1,2,3\}$ and $i<j$, a pair of kernel functions $P_{i}^{3}\oplus P_{j}^{3}(X)$ for $X\in\mathbb{R}_{+}^{n+1}$ is called a biharmonic Poisson kernel associated to a pair of conformal boundary operators $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$, if the following Poisson integral
$v(X)=$ $[P_{i}^{3}\ast f_{i}+P_{j}^{3}\ast f_{j}](X)$
$=$ $\int_{\partial\mathbb{R}_{+}^{n+1}}P_{i}^{3}(x-y,t)f_{i}(y)\mathrm{d}y+\int_{\partial\mathbb{R}_{+}^{n+1}}P_{j}^{3}(x-y,t)f_{j}(y)\mathrm{d}y$
makes sense, provided $f_{i},f_{j}\in C^{2}(\mathbb{R}^{n})$ have suitable decays at infinity, and gives a classical solution to
$\begin{cases}\displaystyle\Delta^{2}v=0&\mathrm{~{}~{}in~{}~{}}\quad~{}~{}\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(v)=f_{i}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{j}^{3}(v)=f_{j}\qquad\,\,\,\,\,\,\,&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (1.6)
Obviously, the notion of biharmonic Poisson kernel could be naturally extended to $\mathbb{B}^{n+1}$ with boundary condition $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$.
Theorem 1.1.
For $n\neq 3$ and $X=(x,t)\in\mathbb{R}_{+}^{n+1}$, we let
$P_{0}^{3}(X)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\frac{t^{3}}{(t^{2}+|x|^{2})^{\frac{n+3}{2}}};$
$P_{1}^{3}(X)=$ $-\frac{2}{|\mathbb{S}^{n}|}\frac{t^{2}}{(t^{2}+|x|^{2})^{\frac{n+1}{2}}};$
$P_{2}^{3}(X)=$ $-\frac{1}{(n-1)|\mathbb{S}^{n}|}\frac{t}{(t^{2}+|x|^{2})^{\frac{n-1}{2}}};$
$P_{3}^{3}(X)=$ $\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\frac{1}{(t^{2}+|x|^{2})^{\frac{n-3}{2}}}.$
Suppose (i) $f_{0}\in C^{2}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ ; (ii) $k\in\{1,2,3\}$, $f_{k}\in C^{2}(\mathbb{R}^{n})$ and $f_{k}=O(|x|^{-k-\delta_{k}})$ as $|x|\to\infty$ for some $\delta_{k}\in\mathbb{R}_{+}$. Then for $i,j\in\{0,1,2,3\}$ and $i+j\neq 3$, the explicit formula of Poisson kernel corresponding to biharmonic boundary value problem (1.6) is $P_{i}^{3}\oplus P_{j}^{3}$.
Remark 1.1.
It is straightforward to verify that other than two pairs $(\mathscr{B}_{0}^{3},\mathscr{B}_{3}^{3})$ and $(\mathscr{B}_{1}^{3},\mathscr{B}_{2}^{3})$, each $(\mathscr{B}_{i}^{3},\mathscr{B}_{j}^{3})$ satisfies ‘the complementing condition’ introduced by Agmon-Douglis-Nirenberg [1, Section 10], see also [15, Section 2.3].
For a smooth compact Riemannian manifold $(X^{n+1},g)$ with boundary $M^{n}$ and $n\geq 4$, under the assumption that a certain conformal invariant $\lambda_{1}(P_{4}^{g})>0$ (see [7]), the existence and uniqueness of smooth solutions to the following boundary value problem was studied in [7, Proposition 1.3]: For $\bar{f}_{0},\bar{f}_{1}\in C^{\infty}(M^{n})$,
$\begin{cases}\displaystyle P_{4}^{g}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad X^{n+1},\\ \displaystyle\mathscr{B}^{3}_{0}U=\bar{f}_{0}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,M^{n},\\ \displaystyle\mathscr{B}^{3}_{1}U=\bar{f}_{1}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,M^{n}.\end{cases}$
On the model space $(\mathbb{B}^{n+1},\mathbb{S}^{n},|\mathrm{d}\xi|^{2})$, for each proper pair of conformal boundary operators, such a biharmonic boundary value problem is expected to be better understood, serving as a first step to the prescribing boundary curvature problem on manifolds with boundary.
Theorem 1.2.
Suppose $n\neq 3$ and $\bar{f}_{i}\in C^{\infty}(\mathbb{S}^{n})$ for $i\in\{0,1,2,3\}$. For $\xi\in\mathbb{B}^{n+1}$ and $\eta\in\mathbb{S}^{n}$, we let
$\bar{P}_{0}^{3}(\xi,\eta)=$ $\frac{n+1}{4|\mathbb{S}^{n}|}\frac{\left(1-|\xi|^{2}\right)^{3}}{|\xi-\eta|^{n+3}};$
$\bar{P}_{1}^{3}(\xi,\eta)=$ $-\frac{1}{2|\mathbb{S}^{n}|}\frac{(1-|\xi|^{2})^{2}}{|\xi-\eta|^{n+1}};$
$\bar{P}_{2}^{3}(\xi,\eta)=$ $-\frac{1}{2(n-1)|\mathbb{S}^{n}|}\frac{1-|\xi|^{2}}{|\xi-\eta|^{n-1}};$
$\bar{P}_{3}^{3}(\xi,\eta)=$ $\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\frac{1}{|\xi-\eta|^{n-3}}.$
Then, for $i<j\in\{0,1,2,3\}$ we have
  • (i)
    For $i+j\neq 3$, the solution of the follow biharmonic boundary value problem
    $\begin{cases}\displaystyle\Delta^{2}U=0\qquad\quad&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}U=\bar{f}_{i}\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}U=\bar{f}_{j}\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\end{cases}$ (1.7)
    is unique and has the form of
    $U(\xi)=\int_{\mathbb{S}^{n}}\bar{P}_{i}^{3}(\xi,\eta)\bar{f}_{i}(\eta)\mathrm{d}V_{\mathbb{S}^{n}}(\eta)+\int_{\mathbb{S}^{n}}\bar{P}_{j}^{3}(\xi,\eta)\bar{f}_{j}(\eta)\mathrm{d}V_{\mathbb{S}^{n}}(\eta).$
  • (ii)
    For pairs $(0,3)$ and $(1,2)$, problem (1.7) has multiple smooth solutions.
Remark 1.2.
It is notable that restricting $\xi$ to $\mathbb{S}^{n}$, $2\bar{P}_{3}^{3}(\xi,\eta)$ coincides with the Green function of fractional GJMS operator $P_{3}$ on $\mathbb{S}^{n}$ (cf. the first author and Shi [10, Theorem 1(3)]). As an application, we can recover a special case in Case [7, Theorem 1.4] about the relationship between extrinsic and intrinsic GJMS operators, see Proposition 3.1 and its biharmonic Poisson kernel proof in Appendix A. Moreover, the non-uniqueness in Theorem 1.2(ii) actually arises from the invertibility of fractional GJMS operators $P_{3}$ and $P_{1}$, see [7, 10] for the definition.
We are ready to introduce a biharmonic Green function on $\mathbb{B}^{n+1}$ with each of four proper pairs of conformal boundary operators. Let
$\Gamma(\xi-\eta)=\Gamma(|\xi-\eta|)=\frac{1}{2(n-1)(n-3)|\mathbb{S}^{n}|}|\xi-\eta|^{3-n}\qquad\mathrm{~{}~{}when~{}~{}}n\neq 3$
and
$\Gamma(\xi-\eta)=\Gamma(|\xi-\eta|)=-\frac{1}{4|\mathbb{S}^{3}|}\log|\xi-\eta|\qquad\qquad\mathrm{~{}~{}when~{}~{}}n=3$
denote the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$. We consider an nonhomogeneous biharmonic boundary value problem with vanishing boundary conditions
$\begin{cases}\Delta^{2}U=\bar{f}\qquad&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{B}^{n+1},\\ \mathscr{B}_{i}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n},\\ \mathscr{B}_{j}^{3}U=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n}.\end{cases}$ (1.8)
Remark 1.3.
A clever observation in (4.1) provides a traditional way in understanding biharmonic Poisson kernel: Conformal boundary operators $\mathscr{B}_{i}^{3}$ indeed act as a bridge between the biharmonic Poisson kernel and fundamental solution.
Definition 1.2 (biharmonic Green function).
Let $\bar{f}\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ and $i<j\in\{0,1,2,3\}$ with $i+j\neq 3$. A function $\bar{G}^{(i,j)}:(\xi,\eta)\in\overline{\mathbb{B}^{n+1}}\times\overline{\mathbb{B}^{n+1}}\mapsto\mathbb{R}\cup\{\infty\}$ for (1.8) is called a Green function, if $\bar{G}^{(i,j)}$ satisfies
  • $\xi\mapsto\bar{G}^{(i,j)}(\xi,\eta)-\Gamma(\xi-\eta)\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$ for all $\eta\in\mathbb{B}^{n+1}$;
  • $\Delta_{\xi}^{2}\left(\bar{G}^{(i,j)}(\xi,\eta)-\Gamma(\xi-\eta)\right)=0$ for all $(\xi,\eta)\in\mathbb{B}^{n+1}\times\mathbb{B}^{n+1}$;
  • $\mathscr{B}_{i}^{3}\bar{G}^{(i,j)}=0,\mathscr{B}_{j}^{3}\bar{G}^{(i,j)}=0$ on $\mathbb{S}^{n}$.
Formally, the above Green function enables us to express a unique solution to (1.8) as
$U(\xi)=\bar{G}^{(i,j)}\ast\bar{f}(\xi)=\int_{\mathbb{B}^{n+1}}\bar{G}^{(i,j)}(\xi,\eta)\bar{f}(\eta)\mathrm{d}\eta.$
We first notice that $\bar{G}^{(0,1)}$ coincides with the Boggio formula (see, e.g., Boggio [2] or Gazzola-Grunau-Sweers [15, Lemma 2.27]), next combine this with biharmonic Poisson kernel to derive explicit formulae of Green functions.
Theorem 1.3.
For $n\neq 3$, the representation formula of Green function $\bar{G}^{(i,j)}$ to (1.8) is
$\bar{G}^{(0,1)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right];$
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right];$
$\bar{G}^{(0,2)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
$\bar{G}^{(1,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
where $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{n+1}}$. Moreover, there holds
$0\leq\bar{G}^{(0,1)}(\xi,\eta)\leq\bar{G}^{(0,2)}(\xi,\eta)\leq\bar{G}^{(1,3)}(\xi,\eta)\leq\bar{G}^{(2,3)}(\xi,\eta),\quad\forall~{}\xi\neq\eta\in\overline{\mathbb{B}^{n+1}},$
with equality if and only if either $\xi\in\mathbb{S}^{n}$ or $\eta\in\mathbb{S}^{n}$.
In the critical dimension that $n=3$, we have
Theorem 1.4.
Suppose $i,j\in\{0,1,2,3\}$ and $i<j$. The following are explicit formulae of biharmonic Poisson kernels on $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
  • (1)
    Under the same assumptions as in Theorem 1.1 when $n=3$, then the same formulae of $P_{i}^{3}$ are true except for
    $P_{3}^{3}(X)=-\frac{1}{4|\mathbb{S}^{3}|}\log(t^{2}+|x|^{2})+C,\qquad C\in\mathbb{R}$
    for $X\in\mathbb{R}_{+}^{4}$, and the biharmonic Poisson kernel for (1.6) is $P_{i}^{3}\oplus P_{j}^{3}$ when $i+j\neq 3$.
  • (2)
    Under the same assumptions as in Theorem 1.2 when $n=3$, then the same formulae of $\bar{P}_{i}^{3}$ are true except for
    $\bar{P}_{3}^{3}(\xi,\eta)=-\frac{1}{2|\mathbb{S}^{3}|}\log|\xi-\eta|+C,\qquad C\in\mathbb{R}$
    for $\xi\in\mathbb{B}^{4},\eta\in\mathbb{S}^{3}$. Moreover, we have
    (i) $i=1,2$ and $j=3$, the solution of (1.7) is
    $U(\xi)-\frac{1}{|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}U\mathrm{d}V_{{\mathbb{S}^{3}}}$
    $=$ $\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta),\quad\forall~{}\xi\in\overline{\mathbb{B}^{4}}.$
    (ii) $(i,j)=(0,1)\mathrm{~{}or~{}}(0,2)$, the unique solution of (1.7) is
    $U(\xi)=\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)+\bar{P}^{3}_{j}(\xi-\eta)\bar{f}_{j}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta),\quad\forall~{}\xi\in\overline{\mathbb{B}^{4}}.$
Theorem 1.5.
The representation formula of Green function $\bar{G}^{(i,j)}$ to (1.8) is
$\bar{G}^{(0,1)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right);$
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)-\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right)+C,\quad C\in\mathbb{R};$
$\bar{G}^{(0,2)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|);$
$\bar{G}^{(1,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)+C,\quad C\in\mathbb{R};$
for $\xi,\eta\in\overline{\mathbb{B}^{4}}$, where $\xi^{\ast}=|\xi|^{-2}\xi$ for $0\neq\xi\in\overline{\mathbb{B}^{4}}$.
We shall assemble the biharmonic Poisson kernel part in Section 3 and the biharmonic Green function part in Section 4.
We next turn to prove classification theorems of nonnegative solutions to (1.5) in $\mathbb{R}_{+}^{n+1}$. Before presenting the proof of our results, in Section 2 we first need to study the nontrivial solutions of an associated homogeneous problem to (1.6) under suitable assumptions on the decay rate at infinity of the initial data, and also describe asymptotic behavior of two typical singular integrals originating from the biharmonic Poisson integrals for $\mathbb{R}_{+}^{n+1}$.
For simplicity, we introduce a ‘geometric bubble function’ associated to above biharmonic boundary value problem by
$U_{\varepsilon,x_{0}}(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\qquad(x,t)\in\mathbb{R}_{+}^{n+1},$
and its boundary function by $u_{\varepsilon,x_{0}}(x)=U_{\varepsilon,x_{0}}(x,0)$, where $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$.
For the pair $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$, under appropriate decay rate at infinity of initial data, the intrinsic nature of $\mathscr{B}^{3}_{3}$ boundary condition enables us to determine the explicit representation of $u$ on $\partial\mathbb{R}_{+}^{n+1}$, see Lemma 5.1. Using basic properties of $U_{\varepsilon,x_{0}}$ and biharmonic Poisson kernel, we establish that every nonnegative smooth solution is a sum of a bubble function and a polynomial. When $p_{3}=p_{3}^{\ast}$, a striking feature is that the bubble function (in general, without geometric meaning) does not vanish even if $p_{1}\neq p_{1}^{\ast}$, and is exact $U_{\varepsilon,x_{0}}$ when $p_{1}=p_{1}^{\ast}$; whereas for $p_{1}<p_{1}^{\ast}$, the bubble function vanishes.
Theorem 1.6.
For $n\geq 4$ and $p_{1},p_{3}>0$, let $u$ be a nonnegative smooth solution of
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{1}^{3}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some constant $c>\max\{3/p_{3},1/p_{1}\}$. Then we have
  • (1)
    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{1}>\frac{1}{n-3}$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{2}$ such that
    $u(x,t)=M_{1}(x,t)+c_{2}t^{2},$
    where
    $M_{1}(x,t)=\left[\mathbb{T}_{1}^{3}P_{1}^{3}\ast u_{\varepsilon,x_{0}}^{p_{1}}+\mathbb{T}_{3}^{3}P_{3}^{3}\ast u_{\varepsilon,x_{0}}^{p_{3}^{\ast}}\right](x,t).$
    In particular, if $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{2}t^{2}.$
  • (2)
    If $0<p_{3}<p_{3}^{\ast}$, then $u(x,t)=c_{2}t^{2}$ for every nonnegative constant $c_{2}$.
For a nonlinear boundary conditions involving the exceptional pair $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$, there still exists a surprising classification theorem for problem (1.5) relying not on $p_{1},p_{2}$ but on $p_{2}/p_{1}$.
Theorem 1.7.
For $n\geq 4$ and $p_{1},p_{2}>0$, let $u$ be a nonnegative smooth solution of
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{1}(u)=\mathbb{T}^{3}_{1}u^{p_{1}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ for some constant $c>2/p_{2}$. Then we have
  • (1)
    If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$ and $p_{1}>0$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{3}$ such that
    $u(x,t)=M_{2}(x,t)+c_{3}t^{3},$
    where
    $M_{2}(x,t)=\left[P_{0}^{3}\ast u_{\varepsilon,x_{0}}^{\frac{p_{1}^{\ast}}{p_{1}}}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast u_{\varepsilon,x_{0}}^{p_{2}^{\ast}}\right](x,t).$
    In particular, $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, then
    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{3}t^{3}.$
  • (2)
    If $\frac{p_{2}}{p_{1}}<\frac{n+1}{n-1}$ , then $u(x,t)=c_{3}t^{3}$ for every nonnegative constant $c_{3}$.
The pair $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$ becomes more challenging and $p_{3}$ is still the dominant exponent. Especially, when $p_{i}=p_{i}^{\ast}$ for $i=1,2,3$, the bubble function happens to be the same $U_{\varepsilon,x_{0}}$; meanwhile, the polynomial part varies according to distinct circumstances.
Theorem 1.8.
For $n\geq 4$ and $p_{2},p_{3}>0$, let $u$ be a nonnegative smooth solution of
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{2}(u)=\mathbb{T}^{3}_{2}u^{p_{2}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Assume $u(x,0)=O(|x|^{-c})$ as $|X|\to\infty$ for some constant $c>\max\{2/p_{2},3/p_{3}\}$. Then we have
  • (1)
    If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then there exist some $\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}$ and nonnegative constant $c_{1}$ such that
    $u(x,t)=M_{3}(x,t)+c_{1}t,$
    where
    $M_{3}(x,t)=\left[P_{0}^{3}\ast u_{\varepsilon,x_{0}}+\mathbb{T}_{2}^{3}P_{2}^{3}\ast u_{\varepsilon,x_{0}}^{p_{2}}\right](x,t).$
    In particular, $p_{2}=p_{2}^{\ast}=\frac{n+1}{n-3}$, then
    $u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{1}t.$
  • (2)
    If $0<p_{3}<\frac{n+3}{n-3}$, then $u(x,t)=c_{1}t$ for every nonnegative constant $c_{1}$.
These classification results demonstrate that the dominant impact of $\mathscr{B}^{3}_{i}$ is different, accordingly, $\mathscr{B}^{3}_{1}\preceq\mathscr{B}^{3}_{2}\preceq\mathscr{B}^{3}_{3}$. In Section 5 we shall present the proof of Theorems 1.6-1.8.
The above classification results in $\mathbb{R}_{+}^{n+1}$ enable us to classify nonnegative solutions to biharmonic boundary value problems on $\mathbb{B}^{n+1}$ with one boundary point singularity.
Theorem 1.9.
Suppose $n\geq 4$ and $U\in C^{4}(\overline{\mathbb{B}^{n+1}}\backslash\{-\mathbf{e}_{n+1}\})$ is a nonnegative solution to
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\},\end{cases}$ (1.9)
where $i,j\in\{1,2,3\}$ with $i<j$. Assume that there exist $\alpha\in(0,\frac{n(n-3)}{n+3})$ and $C\in\mathbb{R}_{+}$ such that
$\sup_{\xi\in\mathbb{S}^{n}\backslash\{-\mathbf{e}_{n+1}\}}U(\xi)|\xi+\mathbf{e}_{n+1}|^{\alpha}\leq C.$ (1.10)
Then there exist some $\xi_{0}\in\mathbb{B}^{n+1}$ and nonnegative constant $\bar{c}$ such that
$U(\xi)=\left(\frac{1-|\xi_{0}|^{2}}{|\xi|^{2}|\xi_{0}|^{2}-2\xi_{0}\cdot\xi+1}\right)^{\frac{n-3}{2}}+\bar{c}|\xi+\mathbf{e}_{n+1}|^{3-n}\left(\frac{1-|\xi|^{2}}{|\xi+\mathbf{e}_{n+1}|}\right)^{6-i-j}.$
Remark 1.4.
If $i=1,j=2$, then the assumption of $\alpha$ can be relaxed to $\alpha\in(0,\frac{(n-1)(n-3)}{n+1})$, which ensures that $n-3-\alpha>\frac{2}{p_{2}^{\ast}}$ by virtue of Theorem 1.7.
A direct consequence of Theorem 1.9 is to prove the uniqueness of positive smooth solutions to a certain type of geometric partial differential equations on $\mathbb{B}^{n+1}$.
Corollary 1.1.
Suppose $n\geq 4$ and $U$ is a positive smooth solution in $\overline{\mathbb{B}^{n+1}}$ to
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)=\mathbb{T}^{3}_{i}U^{\frac{n+2i-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}(U)=\mathbb{T}^{3}_{j}U^{\frac{n+2j-3}{n-3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\mathbb{S}^{n},\end{cases}$
where $i<j\in\{1,2,3\}$. Then
$U(\xi)=\left(\frac{1-|\xi_{0}|^{2}}{|\xi|^{2}|\xi_{0}|^{2}-2\xi_{0}\cdot\xi+1}\right)^{\frac{n-3}{2}}$
for some $\xi_{0}\in\mathbb{B}^{n+1}$.
We now pause to comment on the decay assumptions at infinity of initial data that $u(x,0)=O(|x|^{-c})$ as $|x|\to\infty$ in Theorems 1.6-1.8. For critical exponents $p_{i}^{\ast}$, such assumptions are to ensure that involved Poisson integrals are well-defined. As a result, this deduces that $c>\frac{3(n-3)}{n+3}$ in Theorems 1.6 and 1.8; $c>\frac{2(n-3)}{n+1}$ in Theorem 1.7. As will show in the proof of Theorem 1.9, we know $c=n-3-\alpha$, then the assumption (1.10) leads to $\alpha<\frac{n(n-3)}{n+1}$ and $\alpha<\frac{(n-1)(n-3)}{n+1}$, respectively. Another interpretation of assumption of $\alpha$ is as follows: If $\mathscr{B}^{3}_{3}(U)=\frac{n-3}{2}\tilde{T}^{3}_{3}U^{\frac{n+3}{n-3}}\in L^{1}(\mathbb{S}^{n})$ for a smooth conformal metric $\tilde{g}=U^{4/(n-3)}|\mathrm{d}\xi|^{2}$, then near the singular point $-\mathbf{e}_{n+1}$,
$U^{\frac{n+3}{n-3}}(\xi)\lesssim|\xi+\mathbf{e}_{n+1}|^{-\frac{\alpha(n+3)}{n-3}}\in L^{1}(\mathbb{S}^{n}),$
which requires $0<\alpha<\frac{n(n-3)}{n+3}$. Intuitively, $\mathscr{B}^{3}_{3}(U)\in L^{1}(\mathbb{S}^{n})$ means $U\in W^{3,1}(\mathbb{S}^{n})\hookrightarrow W^{2,\frac{n}{n-1}}(\mathbb{S}^{n})$. This roughly gives $\mathscr{B}^{3}_{2}(U)\in L^{\frac{n}{n-1}}(\mathbb{S}^{n})$, and also
$\left(U^{\frac{n+1}{n-3}}(\xi)\right)^{\frac{n}{n-1}}\lesssim|\xi+\mathbf{e}_{n+1}|^{\frac{\alpha n(n+1)}{(n-1)(n-3)}}\in L^{1}(\mathbb{S}^{n}),$
which requires $\alpha<\frac{(n-1)(n-3)}{n+1}$.
In Section 6, we give the proof of Theorem 1.9, and as applications we also establish two sharp geometric inequalities over some $Q$-curvature flat conformal class of $(\mathbb{B}^{n+1},|\mathrm{d}\xi|^{2})$ for $n\geq 4$, including a sharp inequality of isoperimetric type.
Acknowledgments: The authors thank Professor Qianqiao Guo for bringing [17] into their sight.
2 Preliminaries
We give asymptotic behavior of two prototypes of singular integrals, coming from biharmonic Poisson integrals on $\mathbb{R}_{+}^{n+1}$.
Lemma 2.1.
Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ with $f(x)=O(|x|^{-a})$ as $|x|\to\infty$ for some $a\in\mathbb{R}_{+}$. Assume $20<\alpha<n/$ 2 and $a+2\alpha>n$. We define
$v(x,t)=\int_{\mathbb{R}^{n}}\frac{f(y)}{(t^{2}+|x-y|^{2})^{\alpha}}\mathrm{d}y.$
Then there holds
$\left|v(X)\right|\lesssim$$\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}\quad&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases},\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
Proof..
We decompose $\mathbb{R}^{n}=A_{1}\cup A_{2}\cup A_{3}$ with
$A_{1}=$$\left\{y||y|<|x|/2\right\},\qquad\qquad\qquad\qquad A_{2}=\left\{y||x-y|<|x|/2\right\},$
$A_{3}=$$\left\{y||y|\geq|x|/2,|x-y|\geq|x|/2\right\}.$
Case 1. $|X|\gg 1$ and $t>\frac{|x|}{2}$.
There hold $t\gg 1$ and $|X|\sim t$ due to
$1\leq\frac{|X|^{2}}{t^{2}}=\frac{t^{2}+|x|^{2}}{t^{2}}\leq 5$
In $A_{1}$, we have $|x-y|\sim|x|$ and $t^{2}+|x-y|^{2}\sim t^{2}$, hence
$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{1}}|f(y)|\mathrm{d}y\lesssim$$\begin{cases}t^{-2\alpha}|x|^{n-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ t^{-2\alpha}\log(|x|+1)&\mathrm{if~{}~{}}a=n\end{cases}$
$\lesssim$ $\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
In $A_{2}$, we have $|x-y|<|x|/2<t$, $|y|\sim|x|$ and $t^{2}+|x-y|^{2}\sim t^{2}$, then
$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{2}}|f(y)|\mathrm{d}y\lesssim$$\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
In $A_{3}$, we have $|x-y|\leq|y|+|x|\leq 3|y|$ and $|y|\leq|x-y|+|x|\leq 3|x-y|$, whence $|x-y|\sim|y|$ We further decompose $A_{3}=A_{31}\cup A_{32}$ with
$A_{31}=A_{3}\cap\{y||y|>t\},\qquad\qquad A_{32}=A_{3}\cap\{y||y|\leq t\}.$
In $A_{31}$, we have $t^{2}+|x-y|^{2}\sim|y|^{2}$ and
[…]
In $A_{32}$, we have $t^{2}+|x-y|^{2}\sim t^{2}$ and
$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{32}}|f(y)|\mathrm{d}y\lesssim$$\begin{cases}t^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ t^{-2\alpha}\log t&\mathrm{if~{}~{}}a=n\end{cases}$
$\lesssim$ $\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
Hence we combine the estimates above to obtain
$\left|v(x,t)\right|\lesssim\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}\qquad&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
Case 2. $|X|\gg 1$ and $t\leq\frac{|x|}{2}$.
We have $|x|\gg 1$ and $|X|\sim|x|$ due to
$|x|^{2}\leq|X|^{2}=t^{2}+|x|^{2}\leq\frac{5}{4}|x|^{2}.$
In $A_{1}$, we have $\frac{|x|}{2}\leq|x-y|\leq\frac{3|x|}{2}$ and $t^{2}+|x-y|^{2}\sim|x|^{2}$, then,
$|v(x,t)|\lesssim|x|^{-2\alpha}\int_{A_{1}}|f(y)|\mathrm{d}y\lesssim$$\begin{cases}|x|^{-2\alpha}|x|^{n-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |x|^{-2\alpha}\log(|x|+1)&\mathrm{if~{}~{}}a=n\end{cases}$
$\lesssim$ $\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
In $A_{2}$, we have $|x|/2<|y|<3|x|/2$. Further decompose $A_{2}=A_{21}\cup A_{22}$ with
$A_{21}=A_{2}\cap\{y||x-y|>t\}\qquad\mathrm{and}\qquad A_{22}=A_{2}\cap\{y||x-y|\leq t\}.$
For $A_{21}$, we have $t^{2}+|x-y|^{2}\sim|x-y|^{2}$ and
$|v(x,t)|\lesssim\int_{A_{21}}\frac{|f(y)|}{|x-y|^{2\alpha}}\mathrm{d}y\lesssim|x|^{-a}\int_{|x-y|<|x|/2}\frac{1}{|x-y|^{2\alpha}}\mathrm{d}y\lesssim|X|^{n-2\alpha-a}.$
For $A_{22}$, $t^{2}+|x-y|^{2}\sim t^{2}$, we know
$|v(x,t)|\lesssim t^{-2\alpha}\int_{A_{22}}|f(y)|\mathrm{d}y\lesssim t^{-2\alpha}|x|^{-a}\int_{|x-y|\leq t}\mathrm{d}y\lesssim|X|^{n-2\alpha-a}.$
For $A_{3}$, we know $|y|/3<|x-y|<3|y|$, $|y|>|x|/2>t$, whence, $t^{2}+|x-y|^{2}\sim|y|^{2}$. Then,
$|v(x,t)|\lesssim\int_{A_{3}}\frac{|f(y)|}{|y|^{2\alpha}}\mathrm{d}y\lesssim\int_{|y|>|x|/2}|y|^{-(2\alpha+a)}\mathrm{d}y\lesssim|X|^{n-2\alpha-a}.$
Hence, putting these facts above together we have
$\left|v(x,t)\right|\lesssim\begin{cases}|X|^{n-2\alpha-\min\{a,n\}}\qquad&\mathrm{if~{}~{}}a\neq n\\ |X|^{-2\alpha}\log|X|&\mathrm{if~{}~{}}a=n\end{cases}.$
$\Box$
Lemma 2.2.
Suppose $f\in L_{\mathrm{loc}}^{1}(\mathbb{R}^{n})$ with the decay that $f(x)=O(|x|^{-b})$ as $|x|\to\infty$ for some positive constant $b$. We define
$w(x,t)=\int_{\mathbb{R}^{n}}\frac{t^{\beta}}{(t^{2}+|x-y|^{2})^{\frac{n+\beta}{2}}}f(y)\mathrm{d}y,\quad\beta\in\mathbb{R}_{+}.$
Then there holds
$\left|w(x,t)\right|\lesssim$$\begin{cases}|X|^{-\min\{b,n\}}\qquad&\mathrm{if~{}~{}}b\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}b=n\end{cases},\qquad\quad\mathrm{~{}~{}as~{}~{}}\quad|X|\to\infty.$
Since the same trick works for the above second type singular integral, we omit the proof.
For $r\in\mathbb{R}_{+}$, let $B_{r}^{+}=B_{r}(0)\cap\mathbb{R}^{n+1}_{+}$ and $\partial^{+}B_{r}^{+}=\partial B_{r}^{+}\cap\mathbb{R}^{n+1}_{+}$. Let $P_{k}(x)$ denote a polynomial with degree $k$ in the variable $x$, where $k\in\mathbb{N}$. Given $u\in C^{\infty}(\overline{\mathbb{R}^{n+1}_{+}})$, we introduce a Kelvin transform for $\Delta^{2}$ by
$u^{\ast}(X)=|X|^{3-n}u\left(\frac{X}{|X|^{2}}\right),\qquad X\in\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}.$
For our purpose, we need to refine Sun-Xiong’s results [30, Proposition 3.1 and Theorem 3.2] on homogeneous biharmonic boundary value problems.
Lemma 2.3.
Let $u\in C^{\infty}(\overline{\mathbb{R}^{n+1}_{+}})$ be a solution of
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}u(x,0)=0&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u(x,0)=0\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.1)
Suppose (i) $\lim_{|x|\to\infty}u(x,0)=0$; (ii) there exist some $\delta,C\in\mathbb{R}_{+}$ such that
$u(X)\geq-C|X|^{-\delta}\qquad\mathrm{as}\quad|X|\to\infty.$ (2.2)
Then, $u(x,t)=c_{2}t^{2}$ for some nonnegative constant $c_{2}$.
Proof..
It is notable that the nonnegativity of solutions is removed compared with that of Sun-Xiong [30].
The proof consists of three steps.
Step 1. $u^{\ast}\in L^{1}(B_{1}^{+})$.
First, rewrite the boundary conditions as $\mathscr{B}_{1}^{3}u=0,\mathscr{B}_{3}^{3}u=0$, then it follows from conformal covariance that $u^{\ast}$ satisfies the biharmonic boundary value problem (2.1) in $\overline{\mathbb{R}_{+}^{n+1}}\backslash\{0\}$.
Define
$\eta_{\varepsilon}(t)=\begin{cases}\displaystyle\frac{1}{24}(t-\varepsilon)^{4}\qquad&\mathrm{for}\qquad t\geq\varepsilon,\\ \displaystyle 0\qquad&\mathrm{for}\qquad 0\leq t<\varepsilon.\end{cases}$
Notice that $\eta_{\varepsilon}\in C^{3,1}(\mathbb{R}^{n+1}_{+})$ and $u^{*}$ is biharmonic on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$, then integrating by parts gives
$\int_{B_{1}^{+}}u^{*}(X)\Delta^{2}\eta_{\varepsilon}\mathrm{d}X$
$=$ $\int_{\partial^{+}B_{1}^{+}}\left(\Delta u^{*}(X)\frac{\partial}{\partial r}\eta_{\varepsilon}-\frac{\partial}{\partial r}\Delta u^{*}(X)\eta_{\varepsilon}+u^{*}(X)\frac{\partial}{\partial r}\Delta\eta_{\varepsilon}-\frac{\partial}{\partial r}u^{*}(X)\Delta\eta_{\varepsilon}\right)\mathrm{d}\sigma.$
Letting $\varepsilon\to 0$ we obtain
$\int_{B_{1}^{+}}u^{*}(X)\mathrm{d}X\leq C.$ (2.3)
Write $u^{*}(X)=u_{+}^{*}(X)-u_{-}^{*}(X)$. It follows from the assumption (2.2) that $u_{-}^{*}(X)\leq C|X|^{\delta+3-n}$. Using (2.3) we obtain
$\int_{B_{1}^{+}}u_{+}^{*}(X)\mathrm{d}X\leq C+\int_{B_{1}^{+}}u_{-}^{*}(X)\mathrm{d}X\leq C$
and thus $u^{\ast}\in L^{1}(B_{1}^{+})$.
Step 2. $u(X)=o(|X|^{4})$ as $|X|\to\infty$.
For each $r\in\mathbb{R}_{+}$, we set $v_{r}(X)=u^{*}(rX)$, which is biharmonic in +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$. Now we extend $v_{r}$ by an even reflection to $\mathbb{R}^{n+1}\backslash\{0\}$ and apply the local estimates for biharmonic functions in [27, Proposition 4] to obtain
$\|v_{r}\|_{L^{\infty}(B^{+}_{3/4}\backslash B^{+}_{1/2})}\lesssim\|v_{r}\|_{L^{1}(B^{+}_{5/4}\backslash B^{+}_{1/4})}\lesssim r^{-(n+1)}\|u^{*}\|_{L^{1}(B^{+}_{5r/4}\backslash B^{+}_{r/4})}.$
Since $u^{\ast}\in L^{1}(B_{1}^{+})$, we may take $r$ sufficiently small such that
$\|u^{*}\|_{L^{\infty}(B^{+}_{3r/4}\backslash B^{+}_{r/2})}=o(r^{-(n+1)}).$
This in turn implies
$u(X)=o(|X|^{4})\qquad\mathrm{~{}~{}as~{}~{}}|X|\to\infty.$
Step 3. Completion of the proof.
Extend $u$ by an even reflection
$\tilde{u}(x,t)=\begin{cases}\displaystyle u(x,t)\qquad&\mathrm{for}\qquad t\geq 0,\\ \displaystyle u(x,-t)\qquad&\mathrm{for}\qquad t<0.\end{cases}$
Notice that $\partial_{t}\tilde{u}(x,0)=0$ and
$\partial_{t}^{3}u(x,0)=\partial_{t}\Delta u(x,0)-\partial_{t}\overline{\Delta}u(x,0)=0,$
so $\tilde{u}\in C^{4}(\mathbb{R}^{n+1})$ and is biharmonic in $\mathbb{R}^{n+1}$. By [27, Proposition 4] we obtain
$\|\nabla^{4}\tilde{u}\|_{L^{\infty}(B_{R})}\leq\frac{C}{R^{4}}\frac{1}{|B_{2R}|}\int_{B_{2R}}|\tilde{u}|\mathrm{d}X\leq\frac{o(R^{4})}{R^{4}}\to 0\qquad\mathrm{as}\qquad R\to+\infty.$
Thus, we have
$u(x,t)=\sum_{i=0}^{3}t^{i}P_{3-i}(x).$
The homogeneous boundary conditions on $\partial\mathbb{R}_{+}^{n+1}$ give
$0=\partial_{t}u(x,0)=\sum_{i=1}^{3}it^{i-1}P_{3-i}(x)=P_{2}(x)$
and
$0=\partial^{3}_{t}u(x,0)=\sum_{i=1}^{3}i(i-1)(i-2)t^{i-3}P_{3-i}(x)=6P_{0}(x).$
This yields
$u(x,t)=P_{3}(x)+t^{2}P_{1}(x).$
Since $\liminf_{|X|\to+\infty}u(X)\geq 0$ by virtue of (2.2) and $u(x,0)=o(1)$ as $|x|\to\infty$, we conclude that $P_{3}(x)=0$ and $P_{1}(x)=c_{2}$ for some nonnegative constant $c_{2}$. Hence, the desired assertion follows. $\Box$
Remark 2.1.
If we relax the above assumption (2.2) to $u(X)\geq-C|X|^{\delta}$ for some constant $40<\delta<$ 4, which ensures that $u^{\ast}_{-}\in L^{1}(B_{1}^{+})$, then we apply the argument above to show that
$u(x,t)=t^{2}P_{1}(x).$
Lemma 2.4.
Let $u\in C^{\infty}(\overline{\mathbb{R}^{n+1}_{+}})$ be a solution of
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (2.4)
Suppose there exist some $\delta,C\in\mathbb{R}_{+}$ such that
$u(X)\geq-C|X|^{-\delta}\qquad\mathrm{as}\quad|X|\to\infty.$ (2.5)
Then,
$u(x,t)=tP_{2}(x)+c_{3}t^{3},$
where $c_{3}$ is a nonnegative constant and
$\liminf_{|x|\to\infty}P_{2}(x)\geq 0.$
Proof..
Similarly, the boundary conditions become $\mathscr{B}_{0}^{3}u=0,\mathscr{B}_{3}^{3}u=0$ on $\partial\mathbb{R}_{+}^{n+1}$. By conformal covariance, $u^{\ast}$ satisfies the biharmonic boundary value problem (2.4) on +$\overline{\mathbb{R}^{n+1}_{+}}\backslash\{0\}$. Following the same lines as before, we have $u^{\ast}\in L^{1}(B_{1}^{+})$ and $u(x,t)=o(|X|^{4})$ as $|X|\to\infty$. Finally, we extend $u$ by an odd reflection
$\tilde{u}(x,t)=\begin{cases}\displaystyle u(x,t)\qquad&\mathrm{for}\qquad t\geq 0,\\ \displaystyle-u(x,-t)\qquad&\mathrm{for}\qquad t<0.\end{cases}$
A similar argument as in Lemma 2.3 combined with boundary conditions shows
$u(x,t)=tP_{2}(x)+c_{3}t^{3}.$
This together with the decay assumption at infinity yields the desired assertion. $\Box$
Remark 2.2.
If we relax the above assumption (2.5) to $u(X)\geq-C|X|^{\delta}$ as $|X|\to\infty$ for some $30<\delta<$ 3, then the same result in Lemma 2.4 still holds.
3 Biharmonic Poisson kernel
Our purpose is to derive explicit formulae of biharmonic Poisson kernels for $\mathbb{R}_{+}^{n+1}$ and $\mathbb{B}^{n+1}$.
3.1 Noncritical dimension: The upper half-space
We start with the standard Poisson kernel for $-\Delta$ in $\mathbb{R}_{+}^{n+1}$:
$P(x,t)=\frac{2}{|\mathbb{S}^{n}|}\frac{t}{(t^{2}+|x|^{2})^{\frac{n+1}{2}}}.$
Moreover, $P$ enjoys the property that let $f\in C(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$, then (a) the Poisson integral $P\ast f\in C^{\infty}(\mathbb{R}_{+}^{n+1})\cap L^{\infty}(\mathbb{R}_{+}^{n+1})$ and $\Delta(P\ast f)=0$ in $\mathbb{R}_{+}^{n+1}$; (b) for each fixed $x\in\mathbb{R}^{n}$, $\lim_{t\to 0}P\ast f(x,t)=f(x)$.
Lemma 3.1.
Suppose $f\in C^{2}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ and $\alpha>1$, then for every fixed $x\in\mathbb{R}^{n}$,
$\lim_{t\searrow 0}\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{\alpha}}{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha}{2}}}f(y)\mathrm{d}y=0.$
It is not hard to see that the above result is true provided both $f$ and $|\nabla f|$ are bounded.
Proof..
A direct calculation yields
$\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{\alpha}}{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha}{2}}}f(y)\mathrm{d}y$
$=$ $\int_{\mathbb{R}^{n}}\left[\frac{\alpha t^{\alpha-1}}{{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha}{2}}}}-\frac{(n+\alpha)t^{\alpha+1}}{{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha+2}{2}}}}\right]f(y)\mathrm{d}y$
$=$ $\frac{\alpha}{t}\int_{\mathbb{R}^{n}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y-\frac{n+\alpha}{t}\int_{\mathbb{R}^{n}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha+2}{2}}}\mathrm{d}y.$
For the first term, with some $\sigma\in(0,1)$ we have
$\int_{\mathbb{R}^{n}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y=\int_{|y|>t^{-\sigma}}+\int_{|y|\leq t^{-\sigma}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y:=I_{1}+I_{2}.$
For $I_{1}$, it is not hard to see that
$|I_{1}|\lesssim\int_{|y|>t^{-\sigma}}\frac{1}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y\lesssim t^{\sigma\alpha}.$
For $I_{2}$, as $t|y|<t^{1-\sigma}$ for all $|y|<t^{-\sigma}$, the Taylor expansion together with symmetry enables us to estimate
$I_{2}=$ $\int_{|y|\leq t^{-\sigma}}\frac{f(x)-t\nabla f(x)\cdot y+O(t^{2}|y|^{2})}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y$
$=$ $f(x)\int_{\mathbb{R}^{n}}\frac{1}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y-f(x)\int_{|y|>t^{-\sigma}}\frac{1}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y+\int_{|y|\leq t^{-\sigma}}\frac{O(t^{2}|y|^{2})}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y.$
As $\alpha>1$, we now take $\sigma\in\left(\frac{1}{\alpha},\frac{2}{\alpha}\right)$ for $\alpha\geq 2$ and $\sigma\in(\frac{1}{\alpha},1)$ for $\alpha<2$. Then we have
$I_{2}=\frac{|\mathbb{S}^{n-1}|}{2}B(\frac{n}{2},\frac{\alpha}{2})f(x)+O(t^{\sigma\alpha}).$
Thus, we obtain
$\int_{\mathbb{R}^{n}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha}{2}}}\mathrm{d}y=\frac{|\mathbb{S}^{n-1}|}{2}B(\frac{n}{2},\frac{\alpha}{2})f(x)+O(t^{\sigma\alpha}).$
For the second term, with the same $\sigma$ as above a similar argument yields
$\int_{\mathbb{R}^{n}}\frac{f(x-ty)}{(1+|y|^{2})^{\frac{n+\alpha+2}{2}}}\mathrm{d}y=\frac{|\mathbb{S}^{n-1}|}{2}B(\frac{n}{2},\frac{\alpha+2}{2})f(x)+O(t^{\sigma(\alpha+2)})+O(t^{2}).$
Therefore, with the above choice of $\sigma$ we combine these estimates to conclude that for $t\ll 1$,
$\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{\alpha}}{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+\alpha}{2}}}f(y)\mathrm{d}y$
$=$ $\frac{|\mathbb{S}^{n-1}|}{2t}\left[\alpha B(\frac{n}{2},\frac{\alpha}{2})-(n+\alpha)B(\frac{n}{2},\frac{\alpha+2}{2})\right]f(x)+O(t^{\sigma\alpha-1})$
$=$ $\ O(t^{\sigma\alpha-1}),$
from which the desired conclusion follows. $\Box$
Remark 3.1.
The result in Proposition 3.1 is sharp. Otherwise, if $\alpha=1$ and let $f$ be a nonnegative function in $\mathbb{R}^{n}$ with compact support, then the classical Poisson integral
$u(x,t)=P\ast f(x,t)$
solves $\Delta u=0$ in $\mathbb{R}_{+}^{n+1}$ and $u=f(x)$ on $\partial\mathbb{R}_{+}^{n+1}$. However, for each $x\notin\mathrm{supp}(f)$, the Hopf boundary point lemma yields $\partial_{t}u(x,0)>0$.
Proof of Theorem 1.1. It suffices to show that fix $i,k\in\{0,1,2,3\}$, we define
$[(\mathscr{B}_{i}^{3}P_{k}^{3})\ast f_{k}](x,t)=\int_{\mathbb{R}^{n}}(\mathscr{B}_{i}^{3}P_{k}^{3})(x-y,t)f_{k}(y)\mathrm{d}y,$
then we have
  • (i)
    $\Delta^{2}P_{i}^{3}=0$ in $\mathbb{R}_{+}^{n+1}$.
  • (ii)
    For each pair $(i,k)$ other than $(0,2),(1,3)$, and fixed $x\in\mathbb{R}^{n}$ there hold
    $\lim_{t\to 0}[(\mathscr{B}_{i}^{3}P_{i}^{3})\ast f_{i}](x,t)=f_{i}(x);\qquad\lim_{t\to 0}[(\mathscr{B}_{i}^{3}P_{k}^{3})\ast f_{k}](x,t)=0;$
    $\lim_{t\to 0}[(\mathscr{B}_{k}^{3}P_{k}^{3})\ast f_{k}](x,t)=f_{k}(x);\qquad\lim_{t\to 0}[(\mathscr{B}_{k}^{3}P_{i}^{3})\ast f_{i}](x,t)=0.$
For (i), notice that
$\Delta^{2}|X|^{3-n}=0,\qquad\partial_{t}|X|^{3-n}=(3-n)|X|^{1-n}t,$
this yields that $P_{3}^{3}$ is biharmonic in $\mathbb{R}_{+}^{n+1}$, so is $P_{2}^{3}=\partial_{t}P_{3}^{3}$. Observe that
$\Delta|X|^{1-n}=0,\qquad\partial_{t}|X|^{1-n}=(1-n)|X|^{-1-n}t,$
then
$\Delta^{2}(t\cdot t|X|^{-1-n})=\Delta\left(2\partial_{t}(t|X|^{-1-n})\right)=\frac{2}{1-n}\partial_{t}^{2}\Delta|X|^{1-n}=0,$
from which we know that $P_{1}^{3}$ is biharmonic in $\mathbb{R}_{+}^{n+1}$. Finally, we use
$\partial_{t}(t^{2}|X|^{-1-n})=2t|X|^{-1-n}-(1+n)t^{3}|X|^{-3-n}$ (3.1)
to conclude that $\Delta^{2}P_{0}^{3}=0$ in $\mathbb{R}_{+}^{n+1}$.
In the following, we prove assertion (ii) case by case.
Case 1. Type $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$ boundary condition.
The kernel $P_{0}^{3}$ shares the common property as $P$ that
$\lim_{t\to 0}P_{0}^{3}\ast f_{0}(x,t)=f_{0}(x).$ (3.2)
This is due to
$\lim_{t\to 0}\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y$
$=$ $\lim_{t\to 0}\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{f_{0}(x-ty)}{(1+|y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y=f_{0}(x),$
where the last identity follows from
$\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{1}{(1+|y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y=$ $\frac{2(n+1)|\mathbb{S}^{n-1}|}{|\mathbb{S}^{n}|}\int_{0}^{+\infty}\frac{r^{n-1}}{(1+r^{2})^{\frac{n+3}{2}}}\mathrm{d}r$
$=$ $\frac{(n+1)|\mathbb{S}^{n-1}|}{|\mathbb{S}^{n}|}B\left(\frac{n}{2},\frac{3}{2}\right)=1.$
Here, we emphasize $f_{0}\in C(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})$ is enough to ensure (3.2), same as $P$.
Notice that
$\lim_{t\to 0}P_{1}^{3}\ast f_{1}(x,t)=-\lim_{t\to 0}t[P\ast f_{1}](x,t)=0.$
By Lemma 3.1 we have
$\lim_{t\to 0}[(\mathscr{B}_{1}^{3}P_{0}^{3})\ast f_{0}](x,t)=-\frac{2(n+1)}{|\mathbb{S}^{n}|}\lim_{t\to 0}\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{3}}{\left(t^{2}+|x-y|^{2}\right)^{\frac{n+3}{2}}}f_{0}(y)\mathrm{d}y=0.$
Again by (3.1) we obtain
$\lim_{t\to 0}[(\mathscr{B}_{1}^{3}P_{1}^{3})\ast f_{1}](x,t)=-\lim_{t\to 0}[(\partial_{t}P_{1}^{3})\ast f_{1}](x,t)=\lim_{t\to 0}[(2P-P_{0}^{3})\ast f_{1}](x,t)=f_{1}(x).$
Case 2. Type $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ boundary condition.
Rewrite
$\mathscr{B}^{3}_{2}=\partial_{t}^{2}-\overline{\Delta}=\Delta-2\overline{\Delta}.$
Recall that
$P_{0}^{3}(X)=\frac{2(n+1)}{|\mathbb{S}^{n}|}t^{3}|X|^{-3-n},\quad P_{2}^{3}(X)=-\frac{1}{(n-1)|\mathbb{S}^{n}|}t|X|^{1-n}.$
It has been known in (3.2) that
$\lim_{t\to 0}(P_{0}^{3}\ast f_{0})(x,t)=f_{0}(x).$
Also we have
$\lim_{t\to 0}(P_{2}^{3}\ast f_{2})(x,t)=$ $-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}f_{2}(y)\mathrm{d}y$
$=$ $-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}f_{2}(x-ty)}{(1+|y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y=0,$ (3.3)
where the last identity follows from the following estimate: For each fixed $x$, we choose $M=M(x)>2|x|$ sufficiently large such that $f_{2}(x-ty)=O(|x-ty|^{-2-\delta_{2}})$ for all $|y|>M/t$, then
$\left|\int_{\mathbb{R}^{n}}\frac{t^{2}f_{2}(x-ty)}{(1+|y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y\right|\lesssim$ $\int_{|y|<\frac{M}{t}}\frac{t^{2}}{(1+|y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y+\int_{|y|>\frac{M}{t}}\frac{t^{2}}{|y|^{n-1}|x-ty|^{2+\delta_{2}}}\mathrm{d}y$
$\lesssim$ $t+t^{-\delta_{2}}\int_{|y|>\frac{M}{t}}\frac{1}{|y|^{n+1+\delta_{2}}}\mathrm{d}y\lesssim t.$
Taking $\Delta$ to Equation (3.1) to show
$(n+1)\Delta(t^{3}|X|^{-3-n})=-\partial_{t}\Delta(t\cdot t|X|^{-1-n})=-2\partial_{t}^{2}(t|X|^{-1-n})=2\overline{\Delta}(t|X|^{-1-n}).$
Then we obtain
$\mathscr{B}^{3}_{2}P_{0}^{3}=2\overline{\Delta}(P-P_{0}^{3}).$
This implies
$\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P_{0}^{3})\ast f_{0}](x,t)=2\lim_{t\to 0}\overline{\Delta}[(P-P_{0}^{3})\ast f_{0}](x,t)=0.$
Next we combine
$\Delta(t|X|^{1-n})=2\partial_{t}|X|^{1-n}=2(1-n)t|X|^{-1-n}$
and
$-2\overline{\Delta}(t|X|^{1-n})=2t\partial_{t}^{2}(|X|^{1-n})=2(1-n)[t|X|^{-1-n}-(n+1)t^{3}|X|^{-3-n}]$ (3.4)
to show
$\mathscr{B}^{3}_{2}P_{2}^{3}=P+(P-P_{0}^{3}),$
from which we have
$\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P_{2}^{3})\ast f_{2}](x,t)=f_{2}(x).$ (3.5)
Case 3. Type $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$ boundary condition.
A direct computation yields
$\mathscr{B}^{3}_{1}P^{3}_{1}=\partial_{t}\left(\frac{2}{|\mathbb{S}^{n}|}t^{2}|X|^{-1-n}\right)=$ $\frac{2}{|\mathbb{S}^{n}|}\left(\frac{2t}{(t^{2}+|x|^{2})^{\frac{n+1}{2}}}-\frac{(n+1)t^{3}}{(t^{2}+|x|^{2})^{\frac{n+3}{2}}}\right)$
::=:= $2P-P_{0}^{3},$
Then, it is not hard to see that $\lim_{t\to 0}[(\mathscr{B}^{3}_{1}P^{3}_{1})\ast f_{1}](x,t)=f_{1}(x)$.
Since
$\mathscr{B}^{3}_{1}P^{3}_{3}=\frac{1}{(n-1)|\mathbb{S}^{n}|}t|X|^{1-n},$
similar to (3.1) we have
$\lim_{t\to 0}[(\mathscr{B}_{1}^{3}P_{3}^{3})\ast f_{3}](x,t)=$ $\lim_{t\to 0}\int_{\mathbb{R}^{n}}\frac{1}{(n-1)|\mathbb{S}^{n}|}\frac{tf_{3}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$=$ $\lim_{t\to 0}\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}f_{3}(x-ty)}{(1+|y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y=0.$
Recall that
$\mathscr{B}^{3}_{3}=\partial_{t}\Delta+2\overline{\Delta}\partial_{t}$
and
$P_{1}^{3}(X)=-\frac{2}{|\mathbb{S}^{n}|}t^{2}|X|^{-1-n},\quad P_{3}^{3}(X)=\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}|X|^{3-n}.$
A direct computation yields
$\partial_{t}\Delta(t\cdot t|X|^{-1-n})=$ $2\partial_{t}^{2}(t|X|^{-1-n});$
$2\overline{\Delta}\partial_{t}(t\cdot t|X|^{-1-n})=$ $2\overline{\Delta}\left[2t|X|^{-1-n}-(1+n)t^{3}|X|^{-3-n}\right].$
Since $t|X|^{-1-n}$ is harmonic in $\mathbb{R}_{+}^{n+1}$, we have
$\mathscr{B}^{3}_{3}P_{1}^{3}=-2\overline{\Delta}(P-P_{0}^{3}).$
This directly implies
$\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P_{1}^{3})\ast f_{1}](x,t)=-2\lim_{t\to 0}\overline{\Delta}[(P-P_{0}^{3})\ast f_{1}](x,t)=0.$
A direct computation yields
$\partial_{t}\Delta|X|^{3-n}=$ $(3-n)\Delta(t|X|^{1-n})=2(n-1)(n-3)t|X|^{-1-n};$
$2\overline{\Delta}\partial_{t}|X|^{3-n}=$ $2(3-n)\overline{\Delta}(t|X|^{1-n}).$
These together with (3.4) give
$\mathscr{B}^{3}_{3}P_{3}^{3}=P-\overline{\Delta}(P-P_{0}^{3}).$
This directly implies
$\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P_{3}^{3})\ast f_{3}](x,t)=$ $f_{3}(x)-\overline{\Delta}\lim_{t\to 0}[(P-P_{0}^{3})\ast f_{3}](x,t)=f_{3}(x).$ (3.6)
Case 4. Type $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$ boundary condition.
Since we already have (3.5) and (3.6), it remains to show
$\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P_{3}^{3})\ast f_{3}](x,t)=0,\qquad\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P_{2}^{3})\ast f_{2}](x,t)=0.$
To this end, notice that
$\mathscr{B}^{3}_{2}|X|^{3-n}=$ $\Delta|X|^{3-n}-2\overline{\Delta}|X|^{3-n}$
$=$ $2(3-n)|X|^{1-n}-2(3-n)[(n-1)|X|^{-1-n}t^{2}+|X|^{1-n}]$
$=$ $2(n-1)(n-3)t^{2}|X|^{-1-n},$
whence,
$\mathscr{B}^{3}_{2}P_{3}^{3}=\frac{2}{|\mathbb{S}^{n}|}t^{2}|X|^{-1-n}.$
This implies
$\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P_{3}^{3})\ast f_{3}](x,t)=\lim_{t\to 0}[tP\ast f_{3}(x,t)]=0.$
Since
$\mathscr{B}^{3}_{3}(t|X|^{1-n})=$ $\partial_{t}\Delta(t|X|^{1-n})+2\overline{\Delta}\partial_{t}(t|X|^{1-n})$
$=$ $2\partial_{t}^{2}|X|^{1-n}+2\overline{\Delta}|X|^{1-n}+2t\overline{\Delta}\partial_{t}|X|^{1-n}$
$=$ $2(1-n)t^{2}\overline{\Delta}|X|^{-1-n},$
we have
$\mathscr{B}^{3}_{3}P_{2}^{3}=\frac{2}{|\mathbb{S}^{n}|}\overline{\Delta}(t^{2}|X|^{-1-n}).$
This gives
$\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P_{2}^{3})\ast f_{2}](x,t)=\overline{\Delta}\lim_{t\to 0}[tP\ast f_{2}(x,t)]=0.$
$\Box$
3.2 Noncritical dimension: The unit ball
Our purpose is to give explicit representation formulae of biharmonic Poisson kernels for $\mathbb{B}^{n+1}$ with the help of conformal equivalence between $\mathbb{B}^{n+1}$ and $\mathbb{R}_{+}^{n+1}$,
Let $F:(\mathbb{R}^{n+1}_{+},|\mathrm{d}X|^{2})\to(\mathbb{B}^{n+1},|\mathrm{d}\xi|^{2})$
$\xi=F(X)=-\mathbf{e}_{n+1}+\frac{2(X+\mathbf{e}_{n+1})}{|X+\mathbf{e}_{n+1}|^{2}}$ (3.7)
be a conformal map with the property that
$F^{\ast}(|\mathrm{d}\xi|^{2})=\left(\frac{2}{|X+\mathbf{e}_{n+1}|^{2}}\right)^{2}|\mathrm{d}X|^{2}:=U_{0}(X)^{\frac{4}{n-3}}|\mathrm{d}X|^{2}.$
Clearly, $F$ is exact an inversion of the sphere centered at $-\mathbf{e}_{n+1}$ of radius $\sqrt{2}$ and thus $F=F^{-1}$.
The following important but elementary identity will be frequently used:
$2|\xi+\mathbf{e}_{n+1}|^{-1}|\eta+\mathbf{e}_{n+1}|^{-1}|\xi-\eta|=|F^{-1}(\xi)-F^{-1}(\eta)|,\quad\forall~{}\xi,\eta\in\overline{\mathbb{B}^{n+1}}\setminus\{-\mathbf{e}_{n+1}\}.$ (3.8)
This can be verified by a straightforward calculation.
Given $U\in C^{\infty}(\overline{\mathbb{B}^{n+1}})$, we set
$u=U\circ FU_{0}.$ (3.9)
Clearly, $u(X)=O^{(k)}(|X|^{3-n})$ as $|X|\to\infty$, where $O^{(k)}(|X|^{3-n})$ means the $k$-th derivatives of $v$ belong to $O(|X|^{3-n-k})$ for every $k\in\mathbb{N}$. Moreover, by conformal covariance of the Paneitz operator, we have
$\Delta^{2}u=P_{4}^{|\mathrm{d}X|^{2}}(u)=P_{4}^{|\mathrm{d}X|^{2}}(U\circ FU_{0})=$ $U_{0}^{\frac{n+5}{n-3}}P_{4}^{U_{0}^{\frac{4}{n-3}}|\mathrm{d}X|^{2}}(U\circ F)$
$=$ $U_{0}^{\frac{n+5}{n-3}}(P_{4}^{|\mathrm{d}\xi|^{2}}U)\circ F=U_{0}^{\frac{n+5}{n-3}}(\Delta^{2}U)\circ F.$
Similarly, by (1.4) we have
$\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}X|^{2}}(u)=U_{0}^{\frac{n+2k-3}{n-3}}\left[\left(\mathscr{B}_{k}^{3}\right)_{|\mathrm{d}\xi|^{2}}U\right]\circ F\qquad\mathrm{~{}~{}for~{}~{}}\quad 0\leq k\leq 3.$ (3.10)
Proof of Theorem 1.2. For assertion $(i)$, we first handle $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ type boundary condition.
Let $u$ define as in (3.9), equivalently,
$U(\xi)=\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}u\circ F^{-1}(\xi).$
By virtue of (3.10) we set
$f_{0}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n-3}{2}}\bar{f}_{0}\circ F\left((x,0)\right)\quad\mathrm{and}\quad f_{2}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n+1}{2}}\bar{f}_{2}\circ F((x,0)).$
Then for $|X|,|x|\gg 1$, we know
$u(X)=O(|X|^{3-n}),\quad f_{0}(x)=O(|x|^{3-n}),\quad f_{2}(x)=O(|x|^{-1-n}),$
and $u$ solves
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle u(x,0)=f_{0}\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial^{2}_{t}u(x,0)-\overline{\Delta}u(x,0)=f_{2}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (3.11)
As a direct consequence of Remark 2.2, the solution to the above problem (3.11) is unique. This together with Theorem 1.1 gives
$u(X)=\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y.$
If we let
$\xi=$ $F(X)=\left(\frac{2x}{(1+t)^{2}+|x|^{2}},\frac{1-t^{2}-|x|^{2}}{(1+t)^{2}+|x|^{2}}\right),$
$\eta=$ $F((y,0))=\left(\frac{2y}{1+|y|^{2}},\frac{1-|y|^{2}}{1+|y|^{2}}\right),$
then
$t\circ F^{-1}=\frac{1-|\xi|^{2}}{|\xi+\mathbf{e}_{n+1}|^{2}}$
and from (3.8) that
$2|\xi+\mathbf{e}_{n+1}|^{-1}|\eta+\mathbf{e}_{n+1}|^{-1}|\xi-\eta|=|(x,t)-(y,0)|.$
Therefore, putting these facts together we obtain
$U(\xi)=$ $\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{(t\circ F^{-1})^{3}f_{0}\circ F^{-1}(\eta)\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{n}}{((t\circ F^{-1})^{2}+|(x-y)\circ F^{-1}|^{2})^{\frac{n+3}{2}}}\mathrm{d}y$
$-\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{t\circ F^{-1}f_{2}\circ F^{-1}(\eta)\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{n}}{((t\circ F^{-1})^{2}+|(x-y)\circ F^{-1}|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$=$ $\frac{(n+1)\left(1-|\xi|^{2}\right)^{3}}{4|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{0}(\eta)}{|\xi-\eta|^{n+3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)-\frac{(1-|\xi|^{2})}{2(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{2}(\eta)}{|\xi-\eta|^{n-1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta).$
As indicated in the preceding discussion, for other pairs of conformal boundary operators it suffices to show the uniqueness.
The uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{1})$ is trivial. For $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$, the uniqueness follows from a similar argument above combined with Lemma 2.3 or Remark 2.1.
The remaining boundary condition is $(\mathscr{B}^{3}_{2},\mathscr{B}^{3}_{3})$, the uniqueness becomes a little subtle. Suppose $U$ is a smooth solution of
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{2}U=0&\mathrm{on}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{3}U=0&\mathrm{on}\qquad\mathbb{S}^{n},\end{cases}$
and let $u$ define as in (3.9), then $u$ solves
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}^{2}u(x,0)-\overline{\Delta}u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta u(x,0)+2\overline{\Delta}\partial_{t}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$
with $u(X)=O(|X|^{3-n})$ as $|X|\to\infty$.
By Lemma 2.3 and Theorem 1.1, the unique solution for $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{3})$ is
$u(x,t)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y.$
Since $\partial_{t}u(y,0)=O(|y|^{1-n})$ as $|y|\to\infty$ by definition (3.9) of $u$, we have
$\lim_{t\to 0}\frac{2t}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t\partial_{t}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y=0.$
So, by the uniqueness for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{2})$ due to Lemma 2.4 we obtain $u=0$ and $U=0$ as well.
For assertion $(ii)$, the multiplicity of smooth solutions for the remaining two boundary conditions is a direct consequence of the nex Lemma 3.2. $\Box$
Lemma 3.2.
The following homogeneous biharmonic boundary value problems:
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{0}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{3}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.12)
and
$\begin{cases}\displaystyle\Delta^{2}U=0\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1}\\ \displaystyle\mathscr{B}^{3}_{1}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\\ \displaystyle\mathscr{B}^{3}_{2}(U)=0\qquad&\mathrm{on}\qquad\mathbb{S}^{n}\end{cases}$ (3.13)
have nontrivial smooth solutions.
Proof..
For every $\varphi\in C_{\mathrm{c}}^{\infty}(\mathbb{R}^{n})$, we introduce
$u_{\varphi}(x,t)=-\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\varphi(y)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y$
for the boundary condition $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{3})$; and
$u_{\varphi}(x,t)=\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{\varphi(y)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y$
for the boundary condition $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$.
As in the proof of Theorem 1.1, we know that $u=u_{\varphi}$ solves
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial_{t}\Delta u(x,0)+2\overline{\Delta}\partial_{t}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases}$
and
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1}\\ \displaystyle\partial_{t}u(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\\ \displaystyle\partial^{2}_{t}u(x,0)-\overline{\Delta}u(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}\end{cases},$
respectively. Finally, we may obtain nontrivial smooth solutions of (3.12) and (3.13) via
$U_{\varphi}(\xi)=\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}u_{\varphi}\circ F^{-1}(\xi).$
$\Box$
This also reflects the loss of ‘the complementing condition’ for $(\mathscr{B}^{3}_{0},\mathscr{B}^{3}_{3})$ and $(\mathscr{B}^{3}_{1},\mathscr{B}^{3}_{2})$, under which boundary conditions problem (3.12) or (3.13) fails to have the uniqueness of the solution as explained in [1].
3.3 Critical dimension
Motivated by Chen-Shi [10, Theorem 1(1)] we are able to derive explicit formulae of biharmonic Poisson kernels on $\mathbb{R}_{+}^{4}$ and $\mathbb{B}^{4}$.
Proof of Theorem 1.4. Based on the proof of Theorems 1.1 and 1.2, it suffices to point out several significant differences.
(1) Observe that
$\mathscr{B}^{3}_{3}P^{3}_{3}(X)=(\partial_{t}\Delta+2\overline{\Delta}\partial_{t})P^{3}_{3}(X)=\frac{8}{|\mathbb{S}^{3}|}\frac{t^{3}}{(t^{2}+|x|^{2})^{3}}=P^{3}_{0}(X),$
so, $\lim_{t\to 0}[(\mathscr{B}^{3}_{3}P^{3}_{3})\ast f_{3}](x,t)=f_{3}(x)$. Also, we can apply
$\mathscr{B}^{3}_{1}P^{3}_{3}(X)=$ $\frac{t}{2|\mathbb{S}^{3}|(t^{2}+|x|^{2})}=-P_{2}^{3}(X),$
$\mathscr{B}^{3}_{2}P^{3}_{3}(X)=$ $\frac{2}{|\mathbb{S}^{3}|}\frac{t^{2}}{(t^{2}+|x|^{2})^{2}}=-P_{1}^{3}(X)$
to show
$\lim_{t\to 0}[(\mathscr{B}^{3}_{1}P^{3}_{3})\ast f_{3}](x,t)=0\qquad\mathrm{and}\qquad\lim_{t\to 0}[(\mathscr{B}^{3}_{2}P^{3}_{3})\ast f_{3}](x,t)=0.$
(2) Suppose $U$ is a solution of (1.7). In critical dimension four, we introduce
$u(x,t)=U\circ F(x,t)+\log\frac{2}{(1+t)^{2}+|x|^{2}},\qquad\hat{U}_{0}(x,t):=\log\frac{2}{(1+t)^{2}+|x|^{2}}$
and
$f_{0}(x)=$ $\bar{f}_{0}\circ F(x,0)+\hat{U}(x,0),$
$f_{k}(x)=$ $\left[\bar{f}_{k}\circ F(x,0)+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}\right]e^{k\hat{U}_{0}(x,0)},k\in\{1,2,3\}.$
This essentially follows from conformal covariance of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ that
$e^{2u}|\mathrm{d}X|^{2}=F^{\ast}(e^{2U}|\mathrm{d}\xi|^{2}),\quad e^{2\hat{U}_{0}}|\mathrm{d}X|^{2}=F^{\ast}(|\mathrm{d}\xi|^{2})$
and
$f_{k}=(\mathscr{B}_{k}^{3})_{|\mathrm{d}X|^{2}}(u)=$ $(\mathscr{B}_{k}^{3})_{|\mathrm{d}X|^{2}}(U\circ F)+(\mathscr{B}_{k}^{3})_{|\mathrm{d}X|^{2}}(\hat{U}_{0})$
[…]
$=$ $e^{k\hat{U}_{0}}(\mathscr{B}_{k}^{3})_{F^{\ast}(|\mathrm{d}\xi|^{2})}(U\circ F)+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}e^{k\hat{U}_{0}}$
$=$ $e^{k\hat{U}_{0}}[(\mathscr{B}_{3}^{3})_{|\mathrm{d}\xi|^{2}}(U)]\circ F+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}e^{k\hat{U}_{0}}$
$=$ $e^{k\hat{U}_{0}}(\bar{f}_{k}\circ F+(T_{k}^{3})_{g_{\mathbb{S}^{3}}}).$
In particular, $(T_{1}^{3})_{g_{\mathbb{S}^{3}}}=1,(T_{2}^{3})_{g_{\mathbb{S}^{3}}}=2,(T_{3}^{3})_{g_{\mathbb{S}^{3}}}=4$.
Suppose $u$ is a solution of (1.6). By the biharmonic Poisson kernel on $\mathbb{R}_{+}^{4}$ we have
$u(x,t)=$ $\int_{\partial\mathbb{R}_{+}^{4}}[P_{i}^{3}(x-y,t)f_{i}(y)+P_{j}^{3}(x-y,t)f_{j}(y)]\mathrm{d}y$
$=$ $\int_{\partial\mathbb{R}_{+}^{4}}[P_{i}^{3}(x-y,t)e^{i\hat{U}_{0}(y,0)}f_{i}\circ F(y)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}f_{j}\circ F(y)]\mathrm{d}y+I(x,t),$
where
$I(x,t)=\begin{cases}\int_{\partial\mathbb{R}_{+}^{4}}[P_{0}^{3}(x-y,t)\hat{U}_{0}(y,0)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}(T_{j}^{3})_{g_{\mathbb{S}^{3}}}]\mathrm{d}y,\\ \int_{\partial\mathbb{R}_{+}^{4}}[P_{i}^{3}(x-y,t)e^{i\hat{U}_{0}(y,0)}f_{i}\circ F(y)+P_{j}^{3}(x-y,t)e^{j\hat{U}_{0}(y,0)}f_{j}\circ F(y)]\mathrm{d}y\quad\mathrm{~{}~{}if~{}~{}}i\neq 0.\end{cases}$
Then it is not hard to see that $I=\hat{U}_{0}$ when $(i,j)=(0,1),(0,2)$; $I=\hat{U}_{0}+C,C\in\mathbb{R}$ when $(i,j)=(1,3),(2,3)$.
Hence, using variable changes as in the proof of Theorem 1.2 we obtain
$U(\xi)=\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)+\bar{P}^{3}_{j}(\xi-\eta)\bar{f}_{j}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)$
when $(i,j)=(0,1),(0,2)$; and
$U(\xi)=\int_{\mathbb{S}^{3}}\bar{P}^{3}_{i}(\xi-\eta)\bar{f}_{i}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta)+C,$ (3.14)
when $(i,j)=(1,3),(2,3)$.
It remains to determine the constant $C$ in the following way: On $\mathbb{S}^{3}$, a compatible condition of $\bar{f}_{3}$ arises from the fact that $\mathrm{ker}P_{3}=\mathbb{R}$ by virtue of [10, Theorem 1(1)], more explicitly,
$\int_{\mathbb{S}^{3}}\bar{f}_{3}\mathrm{d}V_{\mathbb{S}^{3}}=\int_{\mathbb{S}^{3}}(-\frac{\partial\Delta U}{\partial r}-2\Delta_{\mathbb{S}^{3}}\frac{\partial U}{\partial r}-2\Delta_{\mathbb{S}^{3}}U)\mathrm{d}V_{\mathbb{S}^{3}}=0.$ (3.15)
Restricting $\xi$ to $\mathbb{S}^{3}$ in (3.14) we find
$U(\xi)=-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}+C.$
Integrating the above equation over $\mathbb{S}^{3}$ to give
$\int_{\mathbb{S}^{3}}U\mathrm{d}V_{\mathbb{S}^{3}}=-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta)\int_{\mathbb{S}^{3}}\log|\xi-\eta|\mathrm{d}V_{\mathbb{S}^{3}}(\xi)+C|\mathbb{S}^{3}|=C|\mathbb{S}^{3}|,$
where we have used (3.15) and the fact that $\int_{\mathbb{S}^{3}}\log|\xi-\eta|\mathrm{d}V_{\mathbb{S}^{3}}(\xi),\forall~{}\eta\in\mathbb{S}^{3}$ is constant. This follows that
$C=\frac{1}{|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}U\mathrm{d}V_{{\mathbb{S}^{3}}}.$
Then the desired assertion follows by inserting the above constant into (3.14). $\Box$
It is not hard to see that Theorem 1.4 could give a direct proof of [6, Proposition 2.2, Lemma 2.3]).
We point out that the proof of Theorems 1.1 and 1.2 is still valid for the case of $n=2$. However, we shall not go deep into this, since it has been well studied by authors [11].
Thanks to biharmonic Poisson kernels on $\mathbb{R}^{n+1}$, we can recover a special case in [7, Theorem 1.4], which clarifies the relationship between extrinsic and intrinsic GJMS operators, and leave its proof to Appendix A for clarity. Something about $P_{1}$ and $P_{3}$.
Proposition 3.1 (cf. [7, Theorem 1.4] and [11, Porposition 2.2]).
Suppose $U$ is a biharmonic function in $\mathbb{B}^{n+1}$ for $n\geq 2$, then there hold
$\mathscr{B}^{3}_{3}U=2P_{3}U\qquad\mathrm{and}\qquad\mathscr{B}^{3}_{2}\left(U\right)=2P_{1}(\mathscr{B}^{3}_{1}U)\qquad\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n},$
where $P_{1},P_{3}$ are the fractional GJMS operators with respect to the round metric, see [7, 10].
Proposition 3.1 indicates that a Dirichlet datum is equivalent to a $\mathscr{B}_{3}^{3}$ datum since $P_{3}$ is invertible; a $\mathscr{B}_{1}^{3}$ datum is equivalent to a $\mathscr{B}_{2}^{3}$ datum since $P_{1}$ is invertible. That is exact the reason why we can construct multiple smooth solutions for a biharmonic equation on $\mathbb{B}^{n+1}$ with these two types of boundary conditions.
4 Biharmonic Green function
For $n\geq 2$, we let $X=(x,t)\in\mathbb{R}_{+}^{n+1},Y=(y,s)\in\mathbb{R}_{+}^{n+1}$. We define $\bar{X}=(x,-t)$, and $\xi=F(X),\eta=F(Y)$ with $F$ being the conformal map as in (3.7).
4.1 Noncritical dimension
Recall that the fundamental solution of $\Delta^{2}$ in $\mathbb{R}^{n+1}$ is given by
$\Gamma(X-Y)=\Gamma(|X-Y|)=\frac{1}{2(n-1)(n-3)|\mathbb{S}^{n}|}|X-Y|^{3-n}.$
Proof of Theorem 1.3. We denote by $G^{(i,j)}$ and $\bar{G}^{(i,j)}$ the Green functions in +$\mathbb{R}^{n+1}_{+}$ and $\mathbb{B}^{n+1}$, respectively.
Fix $X\in\mathbb{R}_{+}^{n+1}$, then for all $Y\in\partial\mathbb{R}_{+}^{n+1}$, direct computations (cf. calculations in the proof of Theorem 1.1) show
$\begin{split}\mathscr{B}^{3}_{0}(\Gamma(X-Y))=&\frac{1}{2}P^{3}_{3}(x-y,t);\\ \mathscr{B}^{3}_{1}(\Gamma(X-Y))=&-\frac{t}{2(n-1)|\mathbb{S}^{n}|(|x-y|^{2}+t^{2})^{\frac{n-1}{2}}}=\frac{1}{2}P^{3}_{2}(x-y,t);\\ \mathscr{B}^{3}_{2}(\Gamma(X-Y))=&\frac{1}{|\mathbb{S}^{n}|}\frac{t^{2}}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}=-\frac{1}{2}P^{3}_{1}(x-y,t);\\ \mathscr{B}^{3}_{3}(\Gamma(X-Y))=&-\frac{(n+1)}{|\mathbb{S}^{n}|}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}=-\frac{1}{2}P^{3}_{0}(x-y,t).\end{split}$ (4.1)
The above formulae demonstrate an intimate connection between the Poisson kernel and fundamental function.
By definition of Green functions on both +$\mathbb{R}^{n+1}_{+}$ and $\mathbb{B}^{n+1}$, we invoke the conformal invariance of $P_{4}^{g}$ to show
$\bar{G}^{(i,j)}(\xi,\eta)=G^{(i,j)}(X,Y)\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}$ (4.2)
by setting $u=U\circ FU_{0}$ and $f=\bar{f}\circ FU_{0}^{(n+5)/(n-3)}$. For brevity, we let
$H^{(i,j)}(X,Y)=G^{(i,j)}(X,Y)-\Gamma(X-Y)$
and
$\bar{H}^{(i,j)}(\xi,\eta)=\bar{G}^{(i,j)}(\xi,\eta)-\Gamma(\xi-\eta).$
Then $H^{(i,j)}$ solves
$\begin{cases}\displaystyle\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad\mathbb{R}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{n+1}.\end{cases}$
For $(i,j)=(0,2)$, we find
$H^{(0,2)}(X,Y)=-\Gamma(\bar{X}-Y).$
This gives
$G^{(0,2)}(X,Y)=\Gamma(X-Y)-\Gamma(\bar{X}-Y).$
Let $\xi^{\ast}=|\xi|^{-2}\xi$ whenever $\xi\neq 0$, then it is not hard to see that $F(\bar{X})=\xi^{\ast}$. Using
$|X+\mathbf{e}_{n+1}||\xi+\mathbf{e}_{n+1}|=2,\quad|\bar{X}+\mathbf{e}_{n+1}||\xi^{\ast}+\mathbf{e}_{n+1}|=2,$
we obtain
$|\xi|^{2}=\frac{|\bar{X}+\mathbf{e}_{n+1}|^{2}}{|X+\mathbf{e}_{n+1}|^{2}}=\frac{|\xi+\mathbf{e}_{n+1}|^{2}}{|\xi^{\ast}+\mathbf{e}_{n+1}|^{2}}.$
Also, the following elementary identity
$2|\xi^{*}-\eta|=|\xi^{\ast}+\mathbf{e}_{n+1}||\eta+\mathbf{e}_{n+1}||\bar{X}-Y|$
gives
$2|\xi||\xi^{*}-\eta|=|\xi+\mathbf{e}_{n+1}||\eta+\mathbf{e}_{n+1}||\bar{X}-Y|.$ (4.3)
Consequently, by (4.2), (3.8) and (4.3) we obtain
$\bar{G}^{(0,2)}(\xi,\eta)=$ $\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}G^{(0,2)}(X,Y)$
$=$ $\Gamma(\xi-\eta)-\Gamma(|\xi|\cdot|\xi^{*}-\eta|).$
For $(i,j)=(1,3)$, similarly we have
$G^{(1,3)}(X,Y)=\Gamma(X-Y)+\Gamma(\bar{X}-Y)$
and
$\bar{G}^{(1,3)}(\xi,\eta)=\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|).$
For $(i,j)=(0,1),(2,3)$, the Boggio formula (see, Boggio [2] or Gazzola-Grunau-Sweers [15, Lemma 2.27]) gives
$\bar{G}^{(0,1)}(\xi,\eta)=$ $\frac{1}{4|\mathbb{S}^{n}|}|\xi-\eta|^{3-n}\int_{1}^{\frac{|\xi||\xi^{\ast}-\eta|}{|\xi-\eta|}}(s^{2}-1)s^{-n}\mathrm{d}s$
$=$ $\Gamma(\xi-\eta)+\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right].$
Set
$f_{2}(y)=-\mathscr{B}^{3}_{2}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{1}(x-y,t)$
and
$f_{3}(y)=-\mathscr{B}^{3}_{3}(\Gamma(X-Y))=\frac{1}{2}P^{3}_{0}(x-y,t).$
We apply Theorem 1.1 together with (4.1) to solve
$H^{(2,3)}(X,Y)=$ $\int_{\mathbb{R}^{n}}P^{3}_{2}(y-z,s)f_{2}(z)\mathrm{d}z+\int_{\mathbb{R}^{n}}P^{3}_{3}(y-z,s)f_{3}(z)\mathrm{d}z$
$=$ $\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{n}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z.$
A similar trick gives
$H^{(0,1)}(X,Y)=-H^{(2,3)}(X,Y).$
Hence, by (3.8) and (4.2) we have
$\bar{H}^{(2,3)}(\xi,\eta)=$ $\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}H^{(2,3)}(X,Y)$
$=$ $-\left(\frac{2}{|\eta+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}H^{(0,1)}(X,Y)=-\bar{H}^{(0,1)}(\xi,\eta),$
from which we conclude that
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)-\frac{1}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left[(n-3)\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}-(n-1)\right].$
Until now, we have already obtained biharmonic Green functions on $\mathbb{R}_{+}^{n+1}$ at the same time.
Observe that
$|\xi|^{2}|\xi^{\ast}-\eta|^{2}=|\xi-\eta|^{2}+(1-|\xi|^{2})(1-|\eta|^{2}),\qquad\forall~{}\xi,\eta\in\overline{\mathbb{B}^{n+1}},$
then it follows that
$\bar{G}^{(2,3)}(\xi,\eta)\geq$ $\Gamma(\xi-\eta)+\Gamma(|\xi|\cdot|\xi^{*}-\eta|)$
$=$ $\bar{G}^{(1,3)}(\xi,\eta)\geq\bar{G}^{(0,2)}(\xi,\eta)\geq 0.$
On the other hand, we have
$[\bar{G}^{(0,2)}-\bar{G}^{(0,1)}](\xi,\eta)=\frac{n-3}{2}\Gamma(|\xi|\cdot|\xi^{\ast}-\eta|)\left(1-\frac{|\xi-\eta|^{2}}{|\xi|^{2}\cdot|\xi^{\ast}-\eta|^{2}}\right)\geq 0.$
Consequently, we combine these facts together to conclude the second assertion. $\Box$
4.2 Critical dimension
The fundamental solution for $\Delta^{2}$ on $\mathbb{R}_{+}^{4}$ is given by
$\Gamma(X-Y)=-\frac{1}{4|\mathbb{S}^{3}|}\log|X-Y|.$
Proof of Theorem 1.5. Fix $X\in\mathbb{R}^{4}_{+}$, then for all $Y\in\partial\mathbb{R}^{4}_{+}$, it is straightforward to show
$\mathscr{B}^{3}_{0}(\Gamma(X-Y))=$ $\frac{1}{2}P^{3}_{3}(x-y,t);$
$\mathscr{B}^{3}_{1}(\Gamma(X-Y))=$ $\frac{1}{2}P^{3}_{2}(x-y,t);$
$\mathscr{B}^{3}_{2}(\Gamma(X-Y))=$ $-\frac{1}{2}P^{3}_{1}(x-y,t);$
$\mathscr{B}^{3}_{3}(\Gamma(X-Y))=$ $-\frac{1}{2}P^{3}_{0}(x-y,t).$
We next derive the Green function on $\mathbb{B}^{4}$ associated with the unique smooth solution of
$\begin{cases}\displaystyle\Delta^{2}U=f\qquad\quad&\mathrm{in}\qquad\mathbb{B}^{4},\\ \displaystyle\mathscr{B}^{3}_{i}U=0\qquad&\mathrm{on}\qquad\mathbb{S}^{3},\\ \displaystyle\mathscr{B}^{3}_{j}U=0\qquad&\mathrm{on}\qquad\mathbb{S}^{3},\end{cases}$
where $\bar{f}\in C^{\infty}(\overline{\mathbb{B}^{4}})$. We claim that
$\bar{G}^{(i,j)}(\xi,\eta)=G^{(i,j)}(X,Y).$
To this end, by conformal covariance (1.3) and (1.4) of $P_{4}^{g}$ and $\mathscr{B}_{k}^{3}$ we have
$\Delta^{2}(U\circ F)=e^{4\hat{U}_{0}}P_{4}^{e^{2\hat{U}_{0}}|\mathrm{d}X|^{2}}(U\circ F)=e^{4\hat{U}_{0}}[P_{4}^{|\mathrm{d}\xi|^{2}}(U)]\circ F=e^{4\hat{U}_{0}}\bar{f}\circ F$
and for $k\in\{0,1,2,3\}$,
$(\mathscr{B}_{k}^{3})_{|\mathrm{d}X|^{2}}(U\circ F)=e^{k\hat{U}_{0}}(\mathscr{B}_{k}^{3})_{e^{2\hat{U}_{0}}|\mathrm{d}X|^{2}}(U\circ F)=e^{k\hat{U}_{0}}[(\mathscr{B}_{k}^{3})_{|\mathrm{d}\xi|^{2}}U]\circ F=0.$
Hence, we are able to apply the Green formula on +$\mathbb{R}^{4}_{+}$ for $U\circ F$ to obtain
$U(\xi)=U\circ F(X)=$ $\int_{\mathbb{R}_{+}^{4}}G^{(i,j)}(X,Y)e^{4\hat{U}_{0}(Y)}\bar{f}\circ F(Y)\mathrm{d}Y$
$=$ $\int_{\mathbb{B}^{4}}G^{(i,j)}(F^{-1}(\xi),F^{-1}(\eta))\bar{f}(\eta)\mathrm{d}\eta,$
which in turn gives
$\bar{G}^{(i,j)}(\xi,\eta)=G^{(i,j)}(F^{-1}(\xi),F^{-1}(\eta)).$
As before, we let
$H^{(i,j)}(X,Y)=G^{(i,j)}(X,Y)-\Gamma(X-Y)$
satisfy
$\begin{cases}\displaystyle\Delta^{2}H^{(i,j)}=0\qquad&\mathrm{in}\qquad\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{i}(H^{(i,j)})=-\mathscr{B}^{3}_{i}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{4}_{+},\\ \displaystyle\mathscr{B}^{3}_{j}(H^{(i,j)})=-\mathscr{B}^{3}_{j}(\Gamma)\qquad&\mathrm{on}\qquad\partial\mathbb{R}^{4}_{+}.\end{cases}$
For $(i,j)=(0,2)$, we find
$H^{(0,2)}(X,Y)=-\Gamma(\bar{X}-Y)$
whence,
$G^{(0,2)}(X,Y)=\Gamma(X-Y)-\Gamma(\bar{X}-Y).$
This together with (3.8), (4.3) gives
$\bar{G}^{(0,2)}(\xi,\eta)=\Gamma(\xi-\eta)-\Gamma(|\xi||\xi^{*}-\eta|).$
We introduce
$\bar{G}^{(1,3)}(\xi,\eta)=\Gamma(\xi-\eta)+\Gamma(|\xi||\xi^{*}-\eta|)+C,\quad C\in\mathbb{R}.$
For $(i,j)=(0,1),(2,3)$, again the Boggio formula gives
$\bar{G}^{(0,1)}(\xi,\eta)=$ $\frac{1}{4|\mathbb{S}^{3}|}\int_{1}^{\frac{|\xi||\xi^{*}-\eta|}{|\xi-\eta|}}(s^{2}-1)s^{-3}\mathrm{d}s$
$=$ $\Gamma(\xi-\eta)-\Gamma(|\xi||\xi^{*}-\eta|)+\frac{1}{8|\mathbb{S}^{3}|}\left(\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right).$
Notice that
$H^{(2,3)}(X,Y)=$ $\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{3}}(P^{3}_{3}(y-z,s)+c)P^{3}_{0}(x-z,t)\mathrm{d}z$
$=$ $\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{2}(y-z,s)P^{3}_{1}(x-z,t)\mathrm{d}z+\frac{1}{2}\int_{\mathbb{R}^{3}}P^{3}_{3}(y-z,s)P^{3}_{0}(x-z,t)\mathrm{d}z+\frac{C}{2},$
then
$H^{(2,3)}(X,Y)=-H^{(0,1)}(X,Y)+\frac{C}{2}.$
Consequently, we have
$\bar{G}^{(2,3)}(\xi,\eta)=$ $\Gamma(\xi-\eta)+\Gamma(|\xi||\xi^{*}-\eta|)-\frac{1}{8|\mathbb{S}^{3}|}\left[\frac{|\xi-\eta|^{2}}{|\xi|^{2}|\xi^{*}-\eta|^{2}}-1\right]+C.$
$\Box$
Eventually, combining the biharmonic Poisson kernel and Green function, we are in a position to present a Green formula for an nonhomogeneous biharmonic boundary value problem on $\mathbb{B}^{n+1}$.
Corollary 4.1 (Green formula for a biharmonic boundary value problem).
For $i,j\in\{0,1,2,3\}$ and $i+j\neq 3$, suppose $\bar{f}\in C^{\infty}(\overline{\mathbb{B}^{n+1}}),\bar{f}_{i},\bar{f}_{j}\in C^{\infty}(\mathbb{S}^{n})$. Then the unique solution of
$\begin{cases}\Delta^{2}U=\bar{f}\qquad&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{B}^{n+1},\\ \mathscr{B}_{i}^{3}U=\bar{f}_{i}&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n},\\ \mathscr{B}_{j}^{3}U=\bar{f}_{j}&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{S}^{n},\end{cases}$
has the form of
$U(\xi)=[\bar{G}^{(i,j)}\ast\bar{f}+\bar{P}_{i}^{3}\ast\bar{f}_{i}+\bar{P}_{j}^{3}\ast\bar{f}_{j}](\xi).$
As an application of the Green formula, we are able to obtain the following comparison principles for biharmonic boundary value problems on $\mathbb{B}^{n+1}$.
Proposition 4.1.
Suppose $n\geq 4$ and $U$ is a smooth function in $\overline{\mathbb{B}^{n+1}}$ satisfying
$\begin{cases}\displaystyle\quad\Delta^{2}U\geq 0\qquad&\mathrm{~{}~{}in~{}~{}}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)\geq 0\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{j}(U)\leq 0\qquad&\mathrm{~{}~{}on~{}~{}}\qquad\mathbb{S}^{n},\end{cases}$
for $i\in\{0,3\},j\in\{1,2\}$, then $U\geq 0$ in $\mathbb{B}^{n+1}$.
Proof..
This is a direct consequent of Corollary 4.1, Theorem 1.2 and Theorem 1.3. $\Box$
5 Classification theorems
As above, let $P_{k}(X)$ and $P_{k}(x)$ denote a polynomial with degree $k$ in the variables $X$ and $x$, where $k\in\mathbb{N}$.
5.1 Proof of Theorem 1.6
Lemma 5.1.
Suppose $u$ is a nonnegative smooth function in $\overline{\mathbb{R}_{+}^{n+1}}$ satisfying
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\mathbb{T}^{3}_{3}u^{p_{3}}\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$
for $0<p_{3}\leq p_{3}^{\ast}=\frac{n+3}{n-3}$. Assume
$u(x,0)=O(|x|^{-c}),\qquad\partial_{t}u(x,0)=O(|x|^{-d})\qquad\mathrm{~{}~{}as~{}~{}}|x|\to\infty$
for some constants $c>3/p_{3}$ and $d>1$. Then we have
  • (i)
    If $p_{3}=p_{3}^{\ast}$, then there exist $\varepsilon>0$ and $x_{0}\in\mathbb{R}^{n}$ such that
    $u(x,0)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}.$
  • (ii)
    If $0<p_{3}<p_{3}^{\ast}$, then $u(x,0)=0$.
Proof..
Notice that $f_{1}(x):=-\partial_{t}u(x,0)=O(|x|^{-d})$ and $f_{3}(x):=\mathbb{T}^{3}_{3}u^{p_{3}}(x,0)=O(|x|^{-p_{3}c})$ as $|x|\to\infty$. We let
$v(x,t)=-\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}f_{1}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+\frac{1}{(n-3)(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{f_{3}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}.$
By Lemmas 2.1 and 2.2, as $|X|\to\infty$ we have
$|v(x,t)|\lesssim t\cdot\begin{cases}|X|^{-\min\{d,n\}}&\mathrm{if~{}~{}}d\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}d=n\end{cases}+\begin{cases}|X|^{3-\min\{p_{3}c,n\}}&\mathrm{if~{}~{}}p_{3}c\neq n\\ |X|^{3-n}\log|X|&\mathrm{if~{}~{}}p_{3}c=n\end{cases}\to 0.$ (5.1)
By Theorem 1.1, $w:=u-v$ solves
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Moreover, since $u$ is nonnegative, by (5.1) we have
$w(x,t)\geq-C|X|^{-\delta}\qquad\mathrm{~{}~{}as~{}~{}}|X|\to\infty,$
where $\delta=\min\{\min\{d,n\}-1,\min\{p_{3}c,n\}-3\}/2>0$, and
$w(x,0)=$ $u(x,0)-\frac{1}{(n-3)(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{f_{3}(y)}{|x-y|^{n-3}}\mathrm{d}y$
$=$ $O(|x|^{\frac{3-\min\{p_{3}c,n\}}{2}})+O(|x|^{-c})\to 0,\qquad\quad\mathrm{~{}~{}as~{}~{}}|x|\to\infty.$
Then it follows from Lemma 2.3 that $w(x,t)=c_{2}^{\prime}t^{2}$ for some nonnegative constant $c_{2}^{\prime}$. This means
$u(x,t)=\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y+\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+c_{2}^{\prime}t^{2}.$ (5.2)
Letting $t\to 0$ we obtain
$u(x,0)=\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{|x-y|^{n-3}}\mathrm{d}y.$
The classification theorem (see Chen-Li-Ou [9] and Y. Y. Li [24]) for the integral equation above gives
$u(x,0)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\quad\varepsilon\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}\qquad\qquad\mathrm{~{}~{}for~{}~{}}p_{3}=p_{3}^{\ast}$
and
$u(x,0)=0\qquad\qquad\mathrm{~{}~{}for~{}~{}}0<p_{3}<p_{3}^{\ast}.$
$\Box$
Proof of Theorem 1.6. For $c>\max\{3/p_{3},1/p_{1}\}$, $u(x,0)=O(|x|^{-c})$ and $\partial_{t}u(x,0)=-\frac{n-3}{2}u^{p_{1}}(x,0)=O(|x|^{-p_{1}c})$ as $|x|\to\infty$.
(1) If $p_{3}=p_{3}^{\ast}=\frac{n+3}{n-3}$, then by Lemma 5.1 we have
$u(x,0)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}.$
Now $u$ solves
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{in}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}u(x,0)=-\frac{n-3}{2}\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{(n-3)p_{1}}{2}}&\mathrm{on}\quad\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}u=\frac{(n^{2}-1)(n-3)}{4}\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n+3}{2}}\quad&\mathrm{on}\quad\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.3)
By (5.2) we know
$u(x,t)=M_{1}(x,t)+c_{2}^{\prime}t^{2}.$ (5.4)
where $c_{2}^{\prime}$ is a nonnegative constant and
$M_{1}(x,t)=$ $\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{1}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{n+3}{2}}\mathrm{d}y$
$-\frac{n-3}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{(n-3)p_{1}}{2}}\mathrm{d}y.$
Especially, assume $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$. On one hand, by Lemmas 2.1 and 2.2 we have
$M_{1}(x,t)=O(|X|^{3-n})\qquad\mathrm{~{}~{}as~{}~{}}|X|\to\infty.$ (5.5)
On the other hand, we consider
$w(x,t)=u(x,t)-\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},$
then $w$ satisfies
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\partial_{t}w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\partial_{t}\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.6)
and for $|X|,|x|\gg 1$,
$w(X)\geq-C|X|^{3-n}\qquad\mathrm{and}\qquad w(x,0)=O(|x|^{-\min\{c,n-3\}}).$
Hence, it follows from Lemma 2.3 that the solution of (5.3) has the form of
$u(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}+c_{2}t^{2}.$ (5.7)
Eventually, we combine (5.4), (5.5) and (5.7) to conclude that $c_{2}^{\prime}=c_{2}$.
(2) If $p_{3}<p_{3}^{\ast}$, then again by Lemma 5.1 we obtain $u(x,0)=0$. In this case, $u$ also solves the same homogeneous biharmonic boundary value problem (5.6), thanks to Lemma 2.3, we conclude that $u(x,t)=c_{2}t^{2}$ for every nonnegative constant $c_{2}$. $\Box$
5.2 Proof of Theorem 1.7
Proof of Theorem 1.7. For $cp_{2}>2$, $f_{0}(x)=u(x,0)=O(|x|^{-c})$ and $f_{2}(x)=\mathbb{T}^{3}_{2}u^{p_{2}}=O(|x|^{-p_{2}c})$ as $|x|\to\infty$. Let
$v(x,t)=\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y.$
with
$|v(x,t)|\lesssim\begin{cases}|X|^{2-\min\{p_{2}c,n\}}&\mathrm{if~{}~{}}p_{2}c\neq n\\ |X|^{2-n}\log|X|&\mathrm{if~{}~{}}p_{2}c=n\end{cases}+\begin{cases}|X|^{-\min\{c,n\}}&\mathrm{if~{}~{}}p_{2}c\neq n\\ |X|^{-n}\log|X|&\mathrm{if~{}~{}}p_{2}c=n\end{cases}\to 0$ (5.8)
as $|X|\to\infty$ by virtue of Lemmas 2.1 and 2.2.
By Theorem 1.1, $w:=u-v$ satisfies
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+}.\end{cases}$
Moreover, by (5.8) we have
$w(x,t)\geq-C|X|^{-\delta}\qquad\mathrm{as}\qquad|X|\to\infty,$
where $\delta=\min\left\{\min\{p_{2}c,n\}-2,\min\{c,n\}\right\}/2>0$. By Lemma 2.4 we have
$u(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tu^{p_{2}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$+tP_{2}(x)+c_{3}t^{3},$ (5.9)
where $c_{3}$ is a nonnegative constant.
From Lemma 3.1 that
$\lim_{t\to 0}\partial_{t}\int_{\mathbb{R}^{n}}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}u(y,0)\mathrm{d}y=0,$
and the first boundary condition we obtain
$\frac{n-3}{2}u^{p_{1}}(x,0)=-\partial_{t}u(x,0)=\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y-P_{2}(x).$ (5.10)
Since $u(x,0)=o(1)$ as $|x|\to\infty$, by Lemma 2.1 we have
$P_{2}(x)=o(1)\qquad\mathrm{as}\qquad|x|\to+\infty,$
so $P_{2}(x)=0$.
Write $k(x)=u^{p_{1}}(x,0)$, the integral equation (5.10) becomes
$k(x)=\frac{1}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{k^{\frac{p_{2}}{p_{1}}}(y)}{|x-y|^{n-1}}\mathrm{d}y.$
(1) If $\frac{p_{2}}{p_{1}}=\frac{n+1}{n-1}$, then we apply the classification theorem (see [9] and [24]) to know
$k(x)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-1}{2}},\qquad\varepsilon\in\mathbb{R}_{+},~{}x_{0}\in\mathbb{R}^{n}.$
Meanwhile, the second boundary condition becomes
$\mathscr{B}^{3}_{2}u=\frac{(n-1)(n-3)}{2}\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n+1}{2}}.$
By (5.2) we have
$u(x,t)=M_{2}(x,t)+c_{3}t^{3}$
and
$M_{2}(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{n-1}{2p_{1}}}\mathrm{d}y$
$-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{n+1}{2}}\mathrm{d}y.$
Especially, when $p_{1}=p_{1}^{\ast}=\frac{n-1}{n-3}$, we consider
$w(x,t)=u(x,t)-\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},$
which satisfies
$\begin{cases}\displaystyle\Delta^{2}w=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle w(x,0)=0\qquad\qquad\qquad\qquad\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\Delta w(x,0)=0\qquad\,\,\,\,\,\,\,&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}^{n+1}_{+},\end{cases}$ (5.11)
with
$w(X)\geq-C|X|^{3-n}\qquad\mathrm{as~{}~{}}|X|\to\infty.$
Moreover, by Lemmas 2.1 and 2.2 we have
$M_{2}(x,t)=O(|X|^{-\min\{\frac{n-1}{p_{1}},n\}})+O(|X|^{2-n})\to 0\qquad\mathrm{as~{}~{}}|X|\to\infty.$
Hence, by Lemma 2.4 we obtain
$u(x,t)=\left(\frac{2\varepsilon}{(\varepsilon+t)^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}}+c_{3}t^{3}.$
(2) If $\frac{p_{2}}{p_{1}}<\frac{n-1}{n-3}$, then the uniqueness theorem in Y. Y. Li [24] yields $u(x,0)=0$ in $\mathbb{R}^{n}$, thus $u(x,t)$ satisfies the homogeneous boundary value problem (5.11). This together with Lemma 2.4 yields $u(x,t)=c_{3}t^{3}$ in $\mathbb{R}_{+}^{n+1}$ for every nonnegative constant $c_{3}$. $\Box$
5.3 Proof of Theorem 1.8
Proof of Theorem 1.8. We now consider
$\begin{cases}\displaystyle\Delta^{2}u=0&\mathrm{~{}~{}in~{}~{}}\quad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}_{2}^{3}(u)=\frac{(n-1)(n-3)}{2}u^{p_{2}}\qquad\qquad&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+},\\ \displaystyle\mathscr{B}_{3}^{3}(u)=\frac{(n^{2}-1)(n-3)}{4}u^{p_{3}}&\mathrm{~{}~{}on~{}~{}}\quad\partial\mathbb{R}^{n+1}_{+}.\end{cases}$ (5.12)
As $cp_{2}>2$, a similar argument as in Theorem 1.7 yields
$u(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tu^{p_{2}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$+tP_{2}(x)+c_{3}t^{3},$ (5.13)
where $c_{3}$ is a nonnegative constant and
$\liminf_{|x|\to\infty}P_{2}(x)\geq 0.$ (5.14)
Moreover,
$-\partial_{t}u(x,0)=\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y-P_{2}(x).$
We first claim that
$P_{2}(x)\equiv c\geq 0\qquad\mathrm{and}\qquad c_{3}=0.$
It is known from (5.14) that $\overline{\Delta}P_{2}(x)$ is a nonnegative constant. With $\tilde{c}_{3}:=c_{3}+\frac{1}{2}\overline{\Delta}P_{2}(x)$ we define
$\tilde{u}(x,t)=$ $u(x,t)-P_{2}(x)t-c_{3}t^{3}+\tilde{c}_{3}t^{3}$
$=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tu^{p_{2}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y+\tilde{c}_{3}t^{3}.$ (5.15)
Clearly, $\tilde{u}$ is biharmonic in $\mathbb{R}_{+}^{n+1}$ and $\tilde{u}(x,0)=u(x,0)$. Recall that $\mathscr{B}^{3}_{2}=\partial_{t}^{2}-\overline{\Delta}$ and $\mathscr{B}^{3}_{3}=\partial_{t}\Delta+2\overline{\Delta}\partial_{t}$, then
$\mathscr{B}^{3}_{3}(\tilde{u})=\mathscr{B}^{3}_{3}(u)+6(\tilde{c}_{3}-c_{3})-3\overline{\Delta}P_{2}(x)=\mathscr{B}^{3}_{3}(u)$
and $\mathscr{B}^{3}_{2}(\tilde{u})=\mathscr{B}^{3}_{2}(u)$ on $\partial\mathbb{R}_{+}^{n+1}$. This means that $\tilde{u}$ also satisfies the biharmonic boundary value problem (5.12), though it might change sign.
On the other hand, notice that $\tilde{u}(x,t)\geq-C|X|^{3}$ as $|X|\to\infty$. By Lemma 2.1 as $|x|\to\infty$ we have
$-\partial_{t}\tilde{u}(x,0)=\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{\tilde{u}^{p_{2}}(y,0)}{|x-y|^{n-1}}\mathrm{d}y=$$\begin{cases}O(|x|^{1-\min\{p_{2}c,n\}})&\qquad\mathrm{if~{}~{}}p_{2}c\neq n\\ O(|x|^{1-n}\log|x|)&\qquad\mathrm{if~{}~{}}p_{2}c=n\end{cases}$
$=$ $O(|x|^{-d}),$
where $d=\min\{cp_{2}-1,(n-1)/2\}>1$. For $c>\max\{2/p_{2},3/p_{3}\}$, we apply Remark 2.1 to obtain
$\tilde{u}(x,t)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}\tilde{u}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y+t^{2}P_{1}(x).$ (5.16)
By (5.3) and (5.16), we can apply Lemmas 2.1 and 2.2 to conclude that
$\tilde{c}_{3}t^{3}-t^{2}P_{1}(x)=o(1)\qquad\mathrm{~{}~{}as~{}~{}}|X|\to\infty.$
This implies $\tilde{c}_{3}=0$ and $P_{1}(x)\equiv 0$. Then, (5.16) becomes
$\tilde{u}(x,t)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}\partial_{t}\tilde{u}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y+\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y.$ (5.17)
Keep in mind that
$0=\tilde{c}_{3}=c_{3}+\frac{\overline{\Delta}P_{2}(x)}{2},$
so, $c_{3}=0$ and $\overline{\Delta}P_{2}(x)=0$. This together with (5.14) implies $P_{2}(x)\equiv c_{1}\geq 0$. Thus, we prove the above claim.
Going back to (5.3) and (5.3) we now have
$u(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tu^{p_{2}}(y,0)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$+c_{1}t$ (5.18)
and
$\tilde{u}(x,t)=u(x,t)-c_{1}t\geq-C|X|\qquad\mathrm{as~{}~{}}|X|\to\infty.$ (5.19)
Furthermore, letting $t\to 0$ in (5.17) we arrive at
$u(x,0)=\tilde{u}(x,0)=\frac{n+1}{4|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u^{p_{3}}(y,0)}{|x-y|^{n-3}}\mathrm{d}y.$
(1) If $p_{3}=\frac{n+3}{n-3}$ and $p_{2}>\frac{2}{n-3}$, then the classification theorem in [9] and [24] gives
$u(x,0)=\left(\frac{2\varepsilon}{\varepsilon^{2}+|x-x_{0}|^{2}}\right)^{\frac{n-3}{2}},\quad x_{0}\in\mathbb{R}^{n},\varepsilon\in\mathbb{R}_{+}.$
This together with (5.3) yields
$u(x,t)=M_{3}(x,t)+c_{1}t,$
where
$M_{3}(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{n-3}{2}}\mathrm{d}y$
$-\frac{n-3}{2|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\left(\frac{2\varepsilon}{\varepsilon^{2}+|y-x_{0}|^{2}}\right)^{\frac{(n-3)p_{2}}{2}}\mathrm{d}y.$
In particular, when $p_{2}=p_{2}^{\ast}$, as before, we notice that $(\tilde{u}-U_{\varepsilon,x_{0}})(X)\geq-C|X|$ as $|X|\to\infty$, and apply Remark 2.2 to conclude that $M_{3}=U_{\varepsilon,x_{0}}$ and then
$u(x,t)=U_{\varepsilon,x_{0}}(x,t)+c_{1}t.$
(2) If $p_{3}<\frac{n+3}{n-3}$, then the uniqueness theorem in [24] implies $u(x,0)=0$. This together with (5.3) implies $u(x,t)=c_{1}t$ in $\mathbb{R}_{+}^{n+1}$ for every nonnegative constant $c_{1}$. $\Box$
6 Applications to biharmonic boundary value problems on balls
Although we focus on the conformal geometry nature for biharmonic boundary value problems, we would like to mention that from a physical point of view, the boundary conditions $(\mathscr{B}_{0}^{3},\mathscr{B}_{1}^{3})$ and $(\mathscr{B}_{0}^{3},\mathscr{B}_{2}^{3})$ have a close connection with the clamped plate and hinged plate model (cf. [15]), respectively. However, the study of other boundary conditions is far less developed.
6.1 One boundary point singularity
Proof of Theorem 1.9. It suffices to give a proof for $i=1$ and $j=3$, since the others are the same in the spirit due to Theorems 1.7 and 1.8.
Let $U$ be a nonnegative solution of (1.9). As before, we introduce
$u(X)=\left(\frac{2}{|X+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}U\circ F.$
Then $u$ solves
$\begin{cases}\displaystyle\Delta^{2}u=0\qquad\qquad\qquad\qquad\,&\mathrm{in}\qquad\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(u)=\mathbb{T}^{3}_{1}u^{\frac{n-1}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{j}(u)=\mathbb{T}^{3}_{3}u^{\frac{n+3}{n-3}}&\mathrm{on}\quad\,\,\,\,\partial\mathbb{R}_{+}^{n+1}.\end{cases}$
By definition of the spherical inversion $F$, we know that $|X+\mathbf{e}_{n+1}||\xi+\mathbf{e}_{n+1}|=2$ and $F(\partial\mathbb{R}_{+}^{n+1})=\partial\mathbb{B}^{n+1}$. By the assumption (1.10) for $U$, we obtain
$u(x,0)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,0)\lesssim|x|^{3-n}|X+\mathbf{e}_{n+1}|^{\alpha}\lesssim|x|^{-(n-3-\alpha)},\quad\mathrm{as~{}~{}}|x|\to\infty$
with $n-3-\alpha>\frac{3(n-3)}{n+3}$. By Theorem 1.6, we have
$u(x,t)=U_{x_{0},\varepsilon}(x,t)+c_{2}t^{2}$
for some $x_{0}\in\mathbb{R}^{n},\varepsilon\in\mathbb{R}_{+}$ and nonnegative constant $c_{2}$.
If we let $\xi_{0}=F(x_{0},\varepsilon)\in\mathbb{B}^{n+1}$, then it is not hard to verify that
$F^{\ast}(U_{\xi_{0}}(\xi)^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2})=U_{x_{0},\varepsilon}(X)^{\frac{4}{n-3}}|\mathrm{d}X|^{2},$
where
$U_{\xi_{0}}(\xi)=\left(\frac{1-|\xi_{0}|^{2}}{|\xi|^{2}|\xi_{0}|^{2}-2\xi_{0}\cdot\xi+1}\right)^{\frac{n-3}{2}},\qquad\xi\in\mathbb{B}^{n+1}.$
In other words,
$U_{\xi_{0}}(\xi)=\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}U_{x_{0},\varepsilon}\circ F^{-1}(\xi).$
Using
$t\circ F^{-1}=\frac{1-|\xi|^{2}}{|\xi+\mathbf{e}_{n+1}|^{2}}$
and (6.1) we obtain
$U(\xi)=$ $\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}u\circ F^{-1}(\xi)$
$=$ $U_{\xi_{0}}(\xi)+c_{2}\left(\frac{2}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{\frac{n-3}{2}}\left(\frac{1-|\xi|^{2}}{|\xi+\mathbf{e}_{n+1}|^{2}}\right)^{2}.$
This follows the desired assertion. $\Box$
6.2 Sharp geometric inequalities
On a compact manifold with boundary, the maximizing problem of the isometric ratio over a scalar-flat conformal class was studied in [22, 21]. With comparison principles in Proposition 4.1, we are motivated to study of the maximizers of the isometric ratio and another sharp geometric inequality over some $Q$-curvature flat conformal class on $(\overline{\mathbb{B}^{n+1}},|\mathrm{d}\xi|^{2})$.
More general, for $n\geq 4$ we define
$\mathcal{M}_{1}=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\};$
$\mathcal{M}_{2}=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}\leq 0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}\geq 0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
and
$\mathring{\mathcal{M}}:=$$\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}(T_{2}^{3})_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
$=$ $\{g\in[|\mathrm{d}\xi|^{2}]~{}|~{}Q_{g}=0\mathrm{~{}~{}in~{}~{}}\mathbb{B}^{n+1},~{}h_{g}=0\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}\}$
by virtue of Proposition 3.1.
Lemma 6.1.
Let $g=U^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}$ be a smooth metric in $\overline{\mathbb{B}^{n+1}}$. Suppose $g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}$. Then we have
  • (1)
    there exists a conformal metric $\tilde{g}=\tilde{U}^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ such that $\tilde{U}\geq U$ in $\mathbb{B}^{n+1}$ with $\tilde{U}=U$ on $\mathbb{S}^{n}$;
  • (2)
    there exists a conformal metric $\tilde{g}=\tilde{U}^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ such that $\tilde{U}\geq U$ in $\mathbb{B}^{n+1}$ with $\mathscr{B}^{3}_{3}(U)=\mathscr{B}^{3}_{3}(\tilde{U})$ on $\mathbb{S}^{n}$.
Proof..
For (i), we introduce an auxiliary function by
$\tilde{U}(\xi)=$ $\frac{(n+1)\left(1-|\xi|^{2}\right)^{3}}{4|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{U(\eta)}{|\xi-\eta|^{n+3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta).$
By Theorem 1.2 we have
$\begin{cases}\displaystyle\quad\Delta^{2}U\leq 0=\Delta^{2}\tilde{U}\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\qquad~{}U=\tilde{U}\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{i}(U)\geq 0=\mathscr{B}^{3}_{i}(\tilde{U})\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\end{cases}$
for $i\in\{1,2\}$. Then the desired assertion follows by Proposition 4.1.
For (ii), we consider
$\tilde{U}(\xi)=\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\mathscr{B}^{3}_{3}(U)(\eta)}{|\xi-\eta|^{n-3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta).$
By Theorem 1.2 we have
$\begin{cases}\displaystyle\quad\Delta^{2}U\leq 0=\Delta^{2}\tilde{U}\qquad&\mathrm{in}\qquad\mathbb{B}^{n+1},\\ \displaystyle\mathscr{B}^{3}_{i}(U)\geq 0=\mathscr{B}^{3}_{i}(\tilde{U})\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\\ \displaystyle\mathscr{B}^{3}_{3}(U)=\mathscr{B}^{3}_{3}(\tilde{U})\qquad&\mathrm{on}\qquad\mathbb{S}^{n},\end{cases}$
for $i\in\{1,2\}$. Then the desired assertion follows from Proposition 4.1. $\Box$
In the following, we consider the isoperimetric ratio over $\mathcal{M}_{1}\cup\mathcal{M}_{2}$:
$\sup_{g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}}\frac{|\mathbb{B}^{n+1}|_{g}}{|\mathbb{S}^{n}|_{g}^{\frac{n+1}{n}}}=\sup_{g\in\mathring{\mathcal{M}}}\frac{|\mathbb{B}^{n+1}|_{g}}{|\mathbb{S}^{n}|_{g}^{\frac{n+1}{n}}}$ (6.1)
by virtue of Lemma 6.1. Given each $g=U^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$, notice that
$\int_{\mathbb{S}^{n}}|(T_{3}^{3})_{g}|^{\frac{2n}{n+3}}\mathrm{d}V_{g}=\left(\frac{2}{n-3}\right)^{\frac{2n}{n+3}}\int_{\mathbb{S}^{n}}|\mathscr{B}^{3}_{3}(u)|^{\frac{2n}{n+3}}\mathrm{d}V_{\mathbb{S}^{n}}.$
then it follows from Theorem 1.2 that $(T_{3}^{3})_{g}$ does not vanish identically. So, again by Lemma 6.1 it is reasonable to study
$\sup_{g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}}\frac{|\mathbb{B}^{n+1}|^{\frac{n-3}{2(n+1)}}_{g}}{\|(T_{3}^{3})_{g}\|_{L^{\frac{2n}{n+3}}(\mathbb{S}^{n},g)}}=\sup_{g\in\mathring{\mathcal{M}}}\frac{|\mathbb{B}^{n+1}|^{\frac{n-3}{2(n+1)}}_{g}}{\|(T_{3}^{3})_{g}\|_{L^{\frac{2n}{n+3}}(\mathbb{S}^{n},g)}}.$ (6.2)
Theorem 6.1.
Let $n\geq 4$. Then for all $g\in\mathcal{M}_{1}\cup\mathcal{M}_{2}$, there hold
$|\mathbb{B}^{n+1}|_{g}\leq d_{n}|\mathbb{S}^{n}|_{g}^{\frac{n+1}{n}}\quad\mathrm{~{}~{}and~{}~{}}\quad|\mathbb{B}^{n+1}|^{\frac{n-3}{2(n+1)}}_{g}\leq e_{n}\|(T_{3}^{3})_{g}\|_{L^{\frac{2n}{n+3}}(\mathbb{S}^{n},g)},$
where
$d_{n}=|\mathbb{S}^{n}|^{-\frac{1}{n}}\int_{0}^{1}r^{n}\big{(}1+\frac{n-3}{4}(1-r^{2})\big{)}^{\frac{2(n+1)}{n-3}}\mathrm{d}r$
and
$e_{n}=|\mathbb{S}^{n}|^{-\frac{7n+3}{2n(n+1)}}\frac{4}{(n^{2}-1)(n-3)}\left[\int_{0}^{1}r^{n}\big{(}1+\frac{n-3}{4}(1-r^{2})\big{)}^{\frac{2(n+1)}{n-3}}\mathrm{d}r\right]^{\frac{n-3}{2(n+1)}}.$
Moreover, the above two equalities hold if and only if $g=cU_{a}^{4/(n-3)}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ with $c\in\mathbb{R}_{+}$ and
$U_{a}(\xi)=\Big{(}\frac{1-|a|^{2}}{|a|^{2}|\xi|^{2}-2a\cdot\xi+1}\Big{)}^{\frac{n-3}{2}}+\frac{n-3}{4}(1-|\xi|^{2})\Big{(}\frac{1-|a|^{2}}{|a|^{2}|\xi|^{2}-2a\cdot\xi+1}\Big{)}^{\frac{n-1}{2}}$ (6.3)
for every $a\in\mathbb{B}^{n+1}$.
Proof..
Given a conformal metric $g=U^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$, if we let
$u(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,0),$
then a similar argument in the proof of Theorem 1.9 yields
$\left(\frac{2}{(1+t)^{2}+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,t)=\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}u(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y.$
To be consistent with notation in Gluck [17], we define
$E_{-2,3}(u)(x,t):=\left(\frac{2}{(1+t)^{2}+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,t).$
So the extremal problem (6.1) is equivalent to
$d_{n}:=\sup_{\begin{subarray}{c}0<u\in C^{\infty}(\mathbb{R}^{n})\\ u=O(|x|^{3-n})\mathrm{~{}as~{}}|x|\to\infty\end{subarray}}\frac{\|E_{-2,3}(u)\|_{L^{\frac{2(n+1)}{n-3}}(\mathbb{R}^{n+1}_{+})}}{\|u\|_{L^{\frac{2n}{n-3}}(\mathbb{R}^{n})}}.$ (6.4)
Gluck [17, Theorem 1.1] states that the optimal constant $d_{n}$ for the extremal problem (6.4) is attained by
$c(|x-x_{0}|^{2}+d^{2})^{\frac{3-n}{2}},\qquad c,d\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n},$
equivalently, modulo of a positive constant, the extremal metric on $\mathbb{S}^{n}$ is a conformal metric $u_{a}^{4/(n-3)}g_{\mathbb{S}^{n}}$ induced by some conformal transformation on $\mathbb{S}^{n}$, which can be expressed as
$u_{a}(\xi)=\Big{(}\frac{1-|a|^{2}}{|a|^{2}|-2a\cdot\xi+1}\Big{)}^{\frac{n-3}{2}}\qquad\mathrm{on~{}~{}}\quad\mathbb{S}^{n}$ (6.5)
for every $a\in\mathbb{B}^{n+1}$. Meanwhile, the uniqueness Theorem 1.2 for both $(\mathscr{B}_{0}^{3},\mathscr{B}_{1}^{3})$ and $(\mathscr{B}_{0}^{3},\mathscr{B}_{2}^{3})$ shows that
$U_{a}(\xi)=\frac{(n+1)\left(1-|\xi|^{2}\right)^{3}}{4|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{u_{a}(\eta)}{|\xi-\eta|^{n+3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)$
such that $U_{a}^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$ is an extremal metric with $U_{a}=u_{a}$ on $\mathbb{S}^{n}$. Indeed, the trick in authors [11, Section 6] can be applied to derive an explicit representation formula of
$U_{a}(\xi)=\Big{(}\frac{1-|a|^{2}}{|a|^{2}|\xi|^{2}-2a\cdot\xi+1}\Big{)}^{\frac{n-3}{2}}+\frac{n-3}{4}(1-|\xi|^{2})\Big{(}\frac{1-|a|^{2}}{|a|^{2}|\xi|^{2}-2a\cdot\xi+1}\Big{)}^{\frac{n-1}{2}}.$
Moreover, we prefer to compute the exact value of $d_{n}$ in a geometric way: Take $1+\frac{n-3}{4}(1-|\xi|^{2})$ as an extremal function, and then
$d_{n}=$ $\frac{\int_{\mathbb{B}^{n+1}}\left(1+\frac{n-3}{4}(1-|\xi|^{2})\right)^{\frac{2(n+1)}{n-3}}\mathrm{d}\xi}{|\mathbb{S}^{n}|^{\frac{n+1}{n}}}$
$=$ $|\mathbb{S}^{n}|^{-\frac{1}{n}}\int_{0}^{1}r^{n}\left(1+\frac{n-3}{4}(1-r^{2})\right)^{\frac{2(n+1)}{n-3}}\mathrm{d}r.$
Similarly, given a conformal metric $g=U^{\frac{4}{n-3}}|\mathrm{d}\xi|^{2}\in\mathring{\mathcal{M}}$, if we let
$u(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n+3}{2}}\mathscr{B}^{3}_{3}(U)\circ F(x,0),$ (6.6)
then
$\left(\frac{2}{(1+t)^{2}+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,t)$
$=$ $\frac{1}{(n-3)(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{u(y)}{(t^{2}+|x-y|^{2})^{\frac{n-3}{2}}}\mathrm{d}y.$
To be consistent with the notation in Dou-Zhu [12], we define
$E_{4}(u)(x,t):=\left(\frac{2}{(1+t)^{2}+|x|^{2}}\right)^{\frac{n-3}{2}}U\circ F(x,t).$
So the extremal problem (6.2) is equivalent to
$\sup_{\begin{subarray}{c}0\not\equiv u\in C^{\infty}(\mathbb{R}^{n})\\ u=O(|x|^{3-n})\mathrm{~{}as~{}}|x|\to\infty\end{subarray}}\frac{\|E_{4}(u)\|_{L^{\frac{2(n+1)}{n-3}}(\mathbb{R}^{n+1}_{+})}}{\|u\|_{L^{\frac{2n}{n+3}}(\mathbb{R}^{n})}}.$ (6.7)
Dou-Zhu [12, Theorem 1.4] states that every maximizer for (6.7) has the form of
$c(|x-x_{0}|^{2}+d^{2})^{-\frac{n+3}{2}},\qquad c,d\in\mathbb{R}_{+},x_{0}\in\mathbb{R}^{n}.$
Combining the above with authors [11, Section 6], we conclude from (6.6) that, modulo a positive constant,
$2P_{3}U=\mathscr{B}^{3}_{3}(U)=\frac{n(n-2)(n-4)}{4}u_{a}^{\frac{n+3}{n-3}}\quad\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n},$
where $u_{a}$ is a conformal factor induced by a conformal transformation on $\mathbb{S}^{n}$ as in (6.5). Since $\mathrm{ker}P_{3}=\{0\}$ implies $U=u_{a}$ on $\mathbb{S}^{n}$, the remaining part of proof is identical to that of (1). $\Box$
Appendix A Appendix: Proof of Proposition 3.1
The combination of Theorem 1.2 and [10, Theorem 1(3)] demonstrates that the Green functions of fractional GJMS operators $P_{1}$ and $P_{3}$ are
$G_{P_{1}}(\xi,\eta)=\frac{2}{(n-1)|\mathbb{S}^{n}|}|\xi-\eta|^{1-n},\quad G_{P_{3}}(\xi,\eta)=2\bar{P}_{3}^{3}(\xi,\eta),\qquad\xi,\eta\in\mathbb{S}^{n}.$
Proof of Proposition 3.1. By (3.10) we set $\bar{f}_{i}=\mathscr{B}^{3}_{i}U$ and
$f_{i}(x)=\left(\frac{2}{1+|x|^{2}}\right)^{\frac{n+2i-3}{2}}\bar{f}_{i}\circ F\left((x,0)\right)$
for $i\in\{0,1,2,3\}$.
When $n\neq 3$, by Theorem 1.2 we have
$U(\xi)=-$ $\frac{(1-|\xi|^{2})^{2}}{2|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{1}(\eta)}{|\xi-\eta|^{n+1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)$
$+$ $\frac{1}{(n-1)(n-3)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{3}(\eta)}{|\xi-\eta|^{n-3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta),\qquad\xi\in\mathbb{B}^{n+1}.$
When $n=4$, by Theorem 1.4 we have
$U(\xi)-\frac{1}{|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}U\mathrm{d}V_{{\mathbb{S}^{3}}}$
$=$ $\int_{\mathbb{S}^{3}}\bar{P}^{3}_{1}(\xi-\eta)\bar{f}_{1}(\eta)\mathrm{d}V_{{\mathbb{S}^{3}}}(\eta)-\frac{1}{2|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}\log|\xi-\eta|\bar{f}_{3}(\eta)\mathrm{d}V_{\mathbb{S}^{3}}(\eta),\quad\forall~{}\xi\in\overline{\mathbb{B}^{4}}.$
Using the elementary formula (3.8) and conformal changes we obtain
$\frac{(1-|\xi|^{2})^{2}}{2|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{1}(\eta)}{|\xi-\eta|^{n+1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)=\frac{2}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{2}f_{1}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+1}{2}}}\mathrm{d}y.$
Since $|\xi|\to 1$ means $t\to 0$, we have
$\lim_{|\xi|\to 1}\frac{(1-|\xi|^{2})^{2}}{2|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{1}(\eta)}{|\xi-\eta|^{n+1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)=0.$
Consequently, we arrive at
$\forall~{}\xi\in\mathbb{S}^{n},\quad U(\xi)=\bar{P}_{3}^{3}\ast\bar{f}_{3}(\xi)=\frac{1}{2}G_{P_{3}}\ast\bar{f}_{3}(\xi),\qquad\mathrm{when~{}~{}}n\neq 3$
and
$\forall~{}\xi\in\mathbb{S}^{n},\quad U(\xi)-\frac{1}{|\mathbb{S}^{3}|}\int_{\mathbb{S}^{3}}U\mathrm{d}V_{\mathbb{S}^{3}}=\bar{P}_{3}^{3}\ast\bar{f}_{3}(\xi)=\frac{1}{2}G_{P_{3}}\ast\bar{f}_{3}(\xi),\qquad\mathrm{when~{}~{}}n=3.$
Consequently, we immediately obtain
$2P_{3}U=\bar{f}_{3}=\mathscr{B}^{3}_{3}U\qquad\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}.$
Similarly, again by Theorem 1.2 we have
$U(\xi)=$ $\frac{(n+1)\left(1-|\xi|^{2}\right)^{3}}{4|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{0}(\eta)}{|\xi-\eta|^{n+3}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)$
$-\frac{(1-|\xi|^{2})}{2(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{2}(\eta)}{|\xi-\eta|^{n-1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta),\qquad\xi\in\mathbb{B}^{n+1}.$
Connecting a function $u$ with $U$ as in (3.9) we employ (3.8) and conformal changes to show
$u(x,t)=$ $\frac{2(n+1)}{|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{t^{3}f_{0}(y)}{(t^{2}+|x-y|^{2})^{\frac{n+3}{2}}}\mathrm{d}y-\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{tf_{2}(y)}{(t^{2}+|x-y|^{2})^{\frac{n-1}{2}}}\mathrm{d}y$
$=$ $[P^{3}_{0}\ast f_{0}+P^{3}_{2}\ast f_{2}](x).$
Then it is not hard to see that
$\lim_{t\to 0}\mathscr{B}^{3}_{1}(u)(x,t)=\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{R}^{n}}\frac{f_{2}(y)}{|x-y|^{n-1}}\mathrm{d}y.$
This together with (3.10) and (3.8) in turn implies that $\forall~{}\xi\in\mathbb{S}^{n}$,
$\mathscr{B}^{3}_{1}U(\xi)=$ $\left[\lim_{t\to 0}\left(\frac{(1+t)^{2}+|x|^{2}}{2}\right)^{\frac{n-1}{2}}\mathscr{B}^{3}_{1}(u)\right]\circ F^{-1}(\xi)$
$=$ $\left[\frac{1}{(n-1)|\mathbb{S}^{n}|}\left(\frac{1+|x|^{2}}{2}\right)^{\frac{n-1}{2}}\int_{\mathbb{R}^{n}}\frac{f_{2}(y)}{|x-y|^{n-1}}\mathrm{d}y\right]\circ F^{-1}(\xi)$
$=$ $\frac{1}{(n-1)|\mathbb{S}^{n}|}\int_{\mathbb{S}^{n}}\frac{\bar{f}_{2}(\eta)}{|\xi-\eta|^{n-1}}\mathrm{d}V_{\mathbb{S}^{n}}(\eta)=\frac{1}{2}G_{P_{1}}\ast\bar{f}_{2},$
which leads to
$2P_{1}\mathscr{B}^{3}_{1}U=\mathscr{B}^{3}_{2}U\qquad\mathrm{~{}~{}on~{}~{}}\mathbb{S}^{n}.$
$\Box$